### v2

# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

Abstract.
Half a century ago Manin showed that given a number field $k$ and a rational prime $\ell$, there exists a uniform bound for the order of cyclic $\ell$-power isogenies between two non-CM elliptic curves over $k$. We generalize this to certain $2$-dimensional families of abelian $3$-folds with multiplication by an imaginary quadratic field.
The first author acknowledges support from the I-SITE ULNE grant ANR-16-IDEX-0004. The second author thanks the Simons Foundation for support from the grant 523557.
Dedicated to the memory of Yuri Manin
Introduction
Given a prime number $\ell$ and a number field $k$, Manin showed in [19] that there exists an integer $r=r(\ell,k)$ such that for any non-CM elliptic curve $E$ over $k$, $E[\ell^{r}]\simeq(\mathbf{Z}/\ell^{r}\mathbf{Z})^{2}$ does not contain a $k$-rational line, or equivalently that the image of the reduction modulo $\ell^{r}$ of its $\ell$-adic Galois representation
$\operatorname{Gal}_{k}=\operatorname{Gal}(\bar{k}/k)\longrightarrow\operatorname{Aut}_{\mathbf{Z}/\ell^{r}\mathbf{Z}}\left(E[\ell^{r}]\right)\simeq\operatorname{GL}(2,\mathbf{Z}/\ell^{r}\mathbf{Z})$
is not contained in a Borel subgroup. Manin’s original proof can be greatly simplified using Faltings’ proof of Mordell’s conjecture, which came later. In a series of papers Cadoret and Tamagawa established a definitive result regarding the uniform boundedness of the $\ell$-primary torsion for $1$-dimensional families of abelian varieties. In this paper we prove an analogous statement for certain $2$-dimensional families of abelian $3$-folds which we believe to be the first result over a genuine two-dimensional base.
Henceforth we fix an imaginary quadratic field $M$ of odd fundamental discriminant $D\neq-3$ and denote by $\mathcal{O}_{M}$ its ring of integers. An abelian $3$-fold of Picard type over a field $k$ containing $M$ will always stand for a principally polarized abelian variety over $k$ of dimension $3$ having multiplication by $\mathcal{O}_{M}$ defined over $k$. Its $\ell$-adic Tate module $T_{\ell}A$ is free of rank $3$ over $\mathcal{O}:=\mathbf{Z}_{\ell}\otimes\mathcal{O}_{M}$ endowed with a continuous $\mathcal{O}$-linear action of $\operatorname{Gal}_{k}$. By a line (resp. plane) in $T_{\ell}A$, we would mean a $\mathcal{O}$-submodule of rank $1$ (resp. $2$) which is a direct factor. More generally, given a positive integer $r$, a line (resp. plane) in $A[\ell^{r}]$ will always be assumed be the image, under the natural reduction map, of a line (resp. plane) in $T_{\ell}A$. Finally, by a full flag we would mean a tuple of a line sitting as direct factor in a plane. Lines (resp. planes) will be called $k$-rational if they are stable by $\operatorname{Gal}_{k}$ (but not necessarily point-wise fixed).
Our first main result addresses the semi-stable case.
Theorem A.
Given a number field $k$, a prime number $\ell$ inert in $M$ and a finite set $S$ of places of $M$, there exists an integer $r=r(\ell,k,S)$ such that for any non-CM abelian $3$-fold AA over $k$ of Picard type which is semi-stable outside $S$, $A[\ell^{r}]$ does not contain a full $k$-rational flag.
As in the case of elliptic curves, the conclusion of Theorem A asks the image of the attached Galois representation
$\operatorname{Gal}_{k}\longrightarrow\operatorname{Aut}_{\mathcal{O}/\ell^{r}\mathcal{O}}\left(A[\ell^{r}]\right)\simeq\mathrm{GU}(3,\mathbf{Z}/\ell^{r}\mathbf{Z})$
not to be contained in a Borel subgroup. Also, as in the case of elliptic curves, it is necessary to cast aside the CM abelian varieties, as their $\ell$-adic representations are potentially reducible.
We next show how one can relax the semi-stability assumption by adding a tiny bit of level structure at $D$. Given a prime $v$ of $M$ above some $p\mid D$, the projective $\operatorname{Gal}_{k}$-action on the $\mathbf{F}_{p}$-vector space $A[v]$ yields a homomorphism $\widetilde{\rho}_{A,p}:\operatorname{Gal}_{k}\to\operatorname{PGL}(2,\mathbf{F}_{p})$ (see (15)). Taking quotient by the unique index two subgroup $\operatorname{PSL}(2,\mathbf{F}_{p})$ of $\operatorname{PGL}(2,\mathbf{F}_{p})$ yields a canonical homomorphism $\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$ and we let $\varepsilon_{A,D}=\prod_{p\mid D}\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$.
Theorem B.
Given a number field $k$ containing $M$ and a prime number $\ell$ inert in $M$, there exists an integer $r=r(\ell,k)$ such that for any non-CM abelian $3$-fold AA over $k$ of Picard type and such that $\varepsilon_{A,D}$ is trivial, $A[\ell^{r}]$ does not contain a full $k$-rational flag.
Theorem B is the main result of this paper and implies Theorem A as follows. Let $k^{\prime}$ be the compositum of the (finitely many) quadratic extensions of $k$ which are unramified outside $S$ and the primes dividing $D$. Given any abelian $3$-fold AA as in Theorem A, we claim that $\varepsilon_{A,D}(\operatorname{Gal}_{k^{\prime}})$ is trivial. Indeed, by a theorem of Grothendieck [15, Prop. 3.5] the semi-stability of $A$ at $v\notin S$, $v\nmid D$ implies that the inertia subgroup of $\operatorname{Gal}_{k}$ at $v$ acts unipotently on the $D$-adic Tate module of $A$, in particular its image by $\varepsilon_{A,D}$ is pro-$D$ hence trivial (as $D$ is odd). Therefore the base change of $A$ to $k^{\prime}$ satisfies the additional assumption in Theorem B, implying that Theorem A holds with $r(\ell,k^{\prime})$ from Theorem B.
For an individual abelian variety $A$, the conclusion of Theorem B is a consequence of the Mumford–Tate conjecture which is known for abelian $3$-folds (see §2.3), so the important feature of the result is its uniformity. As abelian $3$-folds of Picard type are parametrized by Shimura surfaces of Picard type, a natural way to proceed would be to show that the $k$-rational points are not Zariski dense in any of their connected components $Y_{\Gamma}$. Let us for the moment consider the simpler situation from our earlier paper [11] where the congruence subgroups $\Gamma$ were neat. Our method there had two principal steps. The first step involved showing the existence of three linearly independent global holomorphic $1$-forms on the toroidal compactification $X_{\Gamma}$. By a theorem of Faltings concerning the associated Albanese variety this implies that the $k$-rational points on $X_{\Gamma}$ are contained in a divisor $Z$, as predicted by a conjecture of Bombieri and Lang as $X_{\Gamma}$ turns out to be of general type. The second step consisted in applying a result of Nadel requiring $\Gamma$ to be neat and the canonical divisor to be big (in his sense) to deduce that any curve $C$ of genus $\leqslant 1$ contained in $X_{\Gamma}$ is in fact contained in the complement of $Y_{\Gamma}$. Consequently, every curve in $Z$ meeting the open surface $Y_{\Gamma}$ must be of genus $\geqslant 2$ thus, by Faltings’ proof of Mordell’s conjecture for curves, $Y_{\Gamma}(k)$ is finite for any number field $k$.
Let us now say a few words about the techniques involved in the proof of Theorem B. As we are led to consider congruence subgroups of Iwahori type $\Gamma_{0}(\ell^{r})$, which are never neat as they have torsion, both steps mentioned above encounter difficulty and we have to resort to new methods. We produce irregularity by constructing explicit endoscopic automorphic forms in certain non-generic representations $\pi$ on the unitary group in $3$ variables. It is here that the index $2$ projective Galois image condition at $D$, suggested to us by Gross, is essential, as otherwise all of the Picard modular surfaces involved would have trivial Albanese and our approach would not apply for global reasons. Making this strategy actually work yet requires to address some delicate representation theoretic questions to which a significant part of the paper is devoted and on which we will elaborate now.
By Rogawski’s theory $\pi$ is an element of an endoscopic Arthur packet parametrized by an anti-cyclotomic (more precisely, conjugate-symplectic) Hecke character $\lambda$ of $M$. Theorem B imposes conditions implying that $\lambda$ must differ from Gross’ minimally ramified ‘canonical’ characters by a finite order character of $M^{1}$ only ramified at $\ell$. The local Arthur packet at $\ell$ contains two representations, a supercuspidal $\pi_{c,\ell}$ and a non-tempered $\pi_{n,\ell}$, both non-generic. The difficulty of finding $\Gamma_{0}(\ell^{r})$-invariants in $\pi_{c,\ell}$ forces us to work with the global $\pi_{n}$, which is automorphic if, and only if, the global root number $W(\lambda^{3})$ equals +1+1. For $D\equiv 3\pmod{8}$ Gross’ canonical characters work and a computation of matrix coefficients performed in §1.4 shows that the resulting $\pi_{n,\ell}$ has invariants even by the hyperspecial maximal compact subgroup.
When $D\equiv 7\pmod{8}$ the canonical characters yield the wrong sign, leading us to consider $\lambda$’s which are tamely ramified at $\ell$ to switch the sign. It remains however to show that the non-tempered representations $\pi_{n,\ell}$ attached to such $\lambda$’s admit $\Gamma_{0}(\ell^{r})$-invariants for some $r$, which is significantly harder than $\Gamma_{1}(\ell^{r})$-invariants. To that end, in §1.5 we devise a more involved argument relying on Jacquet modules and intertwining operators, and which requires precise computations of exponential sums beyond the reach of known estimates. It would be worthwhile exploring questions regarding levels of non-tempered representations in greater generality.
Once the irregularity of $X_{\Gamma}$ has been shown to be at least $3$ and the Bombieri–Lang conjecture established, one has to deal with the possible curves $C$ of genus $\leqslant 1$ contained in $Y_{\Gamma}$. Using the key Lemma 2.1 on the lack of complex reflections, our Picard modular surfaces only admit a finite number of isolated singularities hence, after removing a finite number of points, $C$ is endowed with an abelian family (see §2.2). This allows us apply the results of Cadoret and Tamagawa regarding the uniform boundedness of the Galois action on the Tate module of such $1$-dimensional families. Finally, each of the finitely many non-CM $k$-rational points are dealt with using the Mumford–Tate conjecture for abelian $3$-folds of Picard type recalled in §2.3, and which is not available in higher dimension.
As our Picard modular surfaces $X_{\Gamma}$ have irregularity $\geqslant 3$, the Enriques–Kodaira classification implies that they are either ruled of genus $q$, or elliptic, or else they are of general type. In the last case, which according to Holzapfel [17, §5.4] occurs for all odd $D\notin\{3,7,11,19,23,31,39,47,71\}$, we show that the Bombieri–Lang Conjecture holds, i.e., that the $k$-rational points are not Zariski dense. Investigating small values of $D$, as suggested by Mazur, seems even more interesting. It is established in loc. cit. that for all $D\neq 71$ in the above list the level $1$ Picard modular surfaces are rational and it would be natural to investigate the nature of their degree $2$ Gross covers that we consider. A way to shed light on this question would be to find an explicit $2$-parameter family of abelian $3$-folds of Picard type to which our theorem applies.
It might be worthwhile remarking that we could have also considered the simpler case of the moduli of principally polarized abelian surfaces $A$ over $k$ with multiplication by $\mathcal{O}_{M}$, which will involve ${\rm U}(1,1)$. However, as ${\rm SU}(1,1)\simeq\operatorname{SL}(2)$ this essentially reduces to the modular curve case. On the other hand, if we consider principally polarized abelian surfaces $A$ with real multiplication, then the family is parametrized by a Hilbert modular surface which has trivial ${\rm H}^{1}$, thus our methods, which rely on the Albanese variety, do not lead to an establishment the Bombieri–Lang Conjecture.
1. Levels for endoscopic non-tempered representations of $\operatorname{U}(3)$
It goes back to the work of Casselman that admissible irreducible representations having non-zero Iwahori invariants are exactly those occurring as sub-quotients in parabolic inductions of unramified characters. Whereas the dimension of the invariants by the depth $r$ Iwahori subgroup in the full induced representation grows as $r$ goes to infinity, this might not always be the case for all its sub-quotients, as shown by the example of the trivial representation of $\operatorname{GL}(2)$, realized as a quotient of a unramified principal series representation.
Another challenging question is to determine which sub-quotient of a parabolically induced unramified character picks up the invariants by a given maximal open compact subgroup. Whereas MacDonald’s formula for zonal spherical functions yields an answer in the case of a maximal hyperspecial subgroup, the general case appears to be an open question.
In this section we fully answer those two natural questions in the case of certain non-tempered endoscopic representations of $\operatorname{U}(3)$ attached to a quadratic extension $E/\mathbf{Q}_{p}$. It will be later applied in a global setting to $E=M_{p}$, where $M$ is an imaginary quadratic field in which the prime $p$ does not split.
In this section of our paper we will adopt local notations.
Let $E$ be a quadratic field extension of $\mathbf{Q}_{p}$, $\mathcal{O}$ be its ring of integers, $\mathcal{P}$ its maximal ideal and $\varpi$ a uniformizer. We assume that $E$ is not a ramified extension of $\mathbf{Q}_{2}$ and we fix a generator $\xi$ of its different ideal such that $\bar{\xi}=-\xi$. Denote by $x\mapsto\bar{x}$ the non-trivial element of $\operatorname{Gal}(E/\mathbf{Q}_{p})$. We fix an additive character $\psi:\mathbf{Q}_{p}\to\mathbf{C}^{\times}$ of conductor 0, i.e. $\ker(\psi)=\mathbf{Z}_{p}$, and we consider the additive character $\psi_{E}$ of $E$ defined as $\psi_{E}(z)=\psi(\operatorname{Tr}_{E/\mathbf{Q}_{p}}(z))$.
Let $G$ be the unique quasi-split unitary group in $3$ variables relative to the extension $E/\mathbf{Q}_{p}$. It can be realized as the automorphisms of $E^{3}$ preserving the hermitian pairing
$\langle x,y\rangle=\bar{x}_{1}y_{3}+\bar{x}_{2}y_{2}+\bar{x}_{3}y_{1}.$
The standard Borel $B$ of $G$ is a product of its torus
$T=\left\{\left(\begin{smallmatrix}\bar{\alpha}&&\\ &\beta&\\ &&\alpha^{-1}\end{smallmatrix}\right)\Big|\alpha\in E^{\times},\beta\in E^{1}\right\}$
with its unipotent subgroup
$N=\left\{[z,x]=\left(\begin{smallmatrix}1&-\bar{z}&\xi x-z\bar{z}/2\\ &1&z\\ &&1\end{smallmatrix}\right)\Big|z\in E,x\in\mathbf{Q}_{p}\right\}.$
1.1. The Bruhat-Tits tree of $\operatorname{U}(3)$
As $G$ has rank $1$, its Bruhat-Tits building is a tree. We will first describe its standard apartment. The relative roots of $G$ are obtained by decomposing the adjoint action on the Lie algebra of the maximal $\mathbf{Q}_{p}$-split torus $T_{0}=\left\{\mathrm{diag}(a,1,a^{-1})|a\in\mathbf{Q}_{p}^{\times}\right\}$ of $G$. The positive elements of the associated root system $\Phi$ are $\{\zeta,2\zeta\}$. Let $h:\mathbf{G}_{m}\to T_{0}\subset G$ be the generator of the co-character lattice $X_{\ast}(T_{0})\simeq\mathbf{Z}$ such that $\langle\zeta,h\rangle=1$. Then the co-root sub-lattice is generated by $\zeta^{\vee}=2h$, so that we have the standard normalization $\langle\zeta,\zeta^{\vee}\rangle=2$. According to [32, §1.15] the affine roots are $\{\pm\zeta+\mathbf{Z}\}\cup\{\pm 2\zeta+\mathbf{Z}\}$ if $E$ is unramified, and $\{\pm\zeta+\tfrac{1}{2}\mathbf{Z}\}\cup\{\pm 2\zeta+\mathbf{Z}+\tfrac{1}{2}\}$ if $E$ is ramified; note that $\delta=0$ in loc. cit. as $E$ is not a ramified extension of $\mathbf{Q}_{2}$. The apartment associated to $T_{0}$ is $\mathbf{R}h$ and its walls are the vanishing sets of these (affine) roots, hence they are given by $\tfrac{1}{2}\mathbf{Z}h=\mathbf{Z}h\cup\tfrac{1}{2}\mathbf{Z}h$, resp. $\tfrac{1}{4}\mathbf{Z}h=\left(\tfrac{1}{2}\mathbf{Z}+\tfrac{1}{4}\right)h\cup\tfrac{1}{2}\mathbf{Z}h$, if $E$ is unramified, resp. ramified. A conjugacy class of maximal compact subgroups can be represented by a wall in the standard apartment. By definition, a wall is hyperspecial if for every $\zeta^{\prime}\in\Phi$ there exists an affine root with gradient $\zeta^{\prime}$ vanishing on that wall. Since $\left(\tfrac{1}{2}\mathbf{Z}+\tfrac{1}{4}\right)h\cap\tfrac{1}{2}\mathbf{Z}h=\varnothing$ this only can happen when $E$ is unramified, in which case the hyperspecial walls are $\mathbf{Z}h\cap\tfrac{1}{2}\mathbf{Z}h=\mathbf{Z}h$. All walls are special, as elements of $\Phi$ are rational multiple of one another.
Let us describe the conjugacy classes of maximal compact subgroups in $G$ in terms equivalent classes of $\mathcal{O}$-lattices $\mathcal{L}$ in $E^{3}$ modulo homothety. Recall the definition of the dual lattice
$\mathcal{L}^{\perp}=\operatorname{Hom}_{\mathcal{O}}(\mathcal{L},\mathcal{O})=\{x\in E^{3}|\langle x,\mathcal{L}\rangle\subset\mathcal{O}\}.$
There are two conjugacy classes of maximal compact subgroups in $G$, those which are stabilizers of self-dual lattices, and those which are stabilizers of almost-self-dual lattices (i.e. $\mathcal{L}$ such that $\mathcal{L}\subsetneq\mathcal{L}^{\perp}\subsetneq\varpi^{-1}\mathcal{L}$). We next give an explicit description of the maximal compact subgroups corresponding to the walls of a chamber in the standard apartment.
The standard maximal compact subgroup $K^{\circ}=G(\mathcal{O})$ of $G$ is defined as the stabilizer of the self-dual lattice $\mathcal{L}^{\circ}=\mathcal{O}^{3}$. It is hyperspecial if and only if $E$ is unramified. The reductive quotient $\overline{G}{}^{\circ}$ is given by $\operatorname{U}(3,\mathbf{F}_{p})$ if $E$ is unramified, and by $\mathrm{O}(3,\mathbf{F}_{p})$ if $E$ is ramified.
The other standard maximal compact subgroup $K^{\prime}$ of $G$, defined as the stabilizer of the almost self-dual lattice $\mathcal{L}^{\prime}=\mathcal{O}\oplus\mathcal{O}\oplus\mathcal{P}$, is given by
$K^{\prime}=\left(\begin{smallmatrix}\mathcal{O}&\mathcal{O}&\mathcal{P}^{-1}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}&\mathcal{P}&\mathcal{O}\end{smallmatrix}\right)\cap G.$
One has $\mathcal{L}^{\prime\perp}=\mathcal{P}^{-1}\oplus\mathcal{O}\oplus\mathcal{O}$ and $K^{\prime}$ acts on $\mathcal{L}^{\prime}/\varpi\mathcal{L}^{\prime\perp}\simeq\mathcal{O}/\mathcal{P}$ via its middle coefficient. The reductive quotient $\overline{G}{}^{\prime}$ is isomorphic to $(\operatorname{U}(1,1)\times\operatorname{U}(1))(\mathbf{F}_{p})$ if $E$ is unramified, and to $\pm\mathbf{1}_{2}\cdot\operatorname{SL}(2,\mathbf{F}_{p})\times\{\pm 1\}$, if $E$ is ramified.
The standard Iwahori subgroup of $G$, defined as $I=K^{\circ}\cap K^{\prime}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$, is the stabilizer of a chamber in the standard apartment in the Bruhat-Tits tree of $G$:
[…]
where $\gamma=\left(\begin{smallmatrix}\varpi^{-1}&&\\ &1&\\ &&\varpi\end{smallmatrix}\right)$ and $I_{2,1}=K^{\circ}\cap\gamma K^{\circ}\gamma^{-1}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}^{2}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$.
One has $G\supset K^{\prime}\supset I\supset I_{2,1}\supset I_{2}$, where $I_{r}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ \mathcal{P}^{r}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}^{r}&\mathcal{P}^{r}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$. Finally, we let $K_{1}\subset K^{\circ}$ be the principal congruence subgroup of matrices $\equiv\mathbf{1}_{3}\pmod{\mathcal{P}}$ and $K_{T}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{P}&\mathcal{P}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{P}\\ \mathcal{P}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$.
1.2. Review of $L$-parameters and $A$-packets
For any integer $n\geqslant 1$ there are exactly two (up to isomorphism) $n$-dimensional hermitian spaces over $E$, depending on the image of the discriminant in $\mathbf{Q}_{p}^{\times}/\mathrm{N}_{E/\mathbf{Q}_{p}}(E^{\times})$, and the corresponding unitary groups $\operatorname{U}(n)$ are isomorphic if and only if $n$ is odd. When $n=2$, by analogy with the Archimedean case, we will denote by $\operatorname{U}(1,1)$ the quasi-split form and by $\operatorname{U}(2)$ the compact one.
The $L$-group of $\operatorname{U}(n)$ is given by $\operatorname{GL}(n,\mathbf{C})\rtimes W_{\mathbf{Q}_{p}}$ with the Weil group $W_{\mathbf{Q}_{p}}$ acting on $\operatorname{GL}(n,\mathbf{C})$ through its quotient $\operatorname{Gal}(E/\mathbf{Q}_{p})$ whose non-trivial element sends $g$ to $w_{n}{}^{t}g^{-1}w_{n}^{-1}$, where $w_{n}$ denotes the anti-diagonal matrix $(1,-1,1,\dots,(-1)^{n-1})$. By definition, an $L$-parameter for the quasi-split $\operatorname{U}(n)$ is a homomorphism $W_{\mathbf{Q}_{p}}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(n,\mathbf{C})\rtimes W_{\mathbf{Q}_{p}}$, but as one knows (see [14, §3]) it is equivalent to ask for its restriction
$\phi:W_{E}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(n,\mathbf{C}),$
to be conjugate-orthogonal if $n$ is odd and conjugate-symplectic if $n$ is even. Recall that $\phi$ is conjugate-self-dual if $\overline{\phi}\simeq\phi^{\vee}$, or equivalently, if the induced representation $\operatorname{Ind}_{W_{E}}^{W_{\mathbf{Q}_{p}}}(\phi)$ is self-dual. Furthermore, $\phi$ is conjugate-orthogonal, resp. conjugate-symplectic, if it preserves a non-degenerate symmetric, resp. skew-symmetric, bilinear form. Note that while Schur’s Lemma implies that any irreducible self-dual (or conjugate-self-dual) parameter has a well defined sign, this need not be always the case for reducible parameters.
[…]
The base change $\nu_{E}(z)=\nu(z/\overline{z})$ from $\operatorname{U}(1)$ to $\operatorname{GL}(1)/E$ of a character $\nu$ of $E^{1}$ is conjugate-orthogonal and conversely any conjugate-orthogonal character of $E^{\times}$ is obtained in that way. For $\lambda$ a conjugate-symplectic character of $E^{\times}$, the conjugate-orthogonal representation
$(\lambda\otimes\operatorname{St})\oplus\nu_{E} :W_{E}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(3,\mathbf{C})$
would be of key relevance to us. It yields an $L$-parameter $\phi_{\lambda,\nu}$ of $G$, coming from an $L$-parameter of the (unique) cuspidal endoscopic subgroup $H=\operatorname{U}(1,1)\times\operatorname{U}(1)$ of $G$. The cardinality of the corresponding $L$-packet $\Pi_{L}(\phi_{\lambda,\nu})$ is given by the order of the centralizer of $\phi_{\lambda,\nu}$ (modulo center) which turns out to be $2$. More precisely, $\Pi_{L}(\phi_{\lambda,\nu})$ contains two discrete series representations $\pi_{2}$ and $\pi_{c}$ of $\operatorname{U}(3)$, exactly one of them, namely $\pi_{c}$, being supercuspidal (see [26, Chap. 12.2] where this $L$-packet is denoted $\Pi_{L}(\operatorname{St}_{H}(\xi))$). There is another endoscopic $L$-packet for $G$ consisting of a single non-tempered representation $\pi_{n}$ whose the $L$-parameter is given by
$\lambda|\cdot|_{E}^{1/2}\oplus\lambda|\cdot|_{E}^{-1/2}\oplus\nu_{E}:W_{E}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(3,\mathbf{C}).$
Rogawski’s theory [26, 27] describes the automorphic representations contributing to the $\operatorname{H}^{1}$ of Shimura surfaces of Picard type in terms global Arthur packets (see [11, §3.1] for a summary). The corresponding local Arthur packet at $p$ has $2$ elements $\Pi(\lambda,\nu)=\{\pi_{n},\pi_{c}\}$ (see [26, §12.3.3], where $\pi_{c}$ is denoted $\pi^{s}$), and the restriction to $W_{E}$ of its $A$-parameter is given by
$(\lambda\otimes\mathbf{1}\otimes\operatorname{St})\oplus\nu_{E}:W_{E}\times\operatorname{SL}(2,\mathbf{C})\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(3,\mathbf{C}),$
while the $A$-parameter of $\pi_{2}$ is given by $(\lambda\otimes\operatorname{St}\otimes\mathbf{1})\oplus\nu_{E}$.
Crucial for us would be the description $\pi_{n}$ and $\pi_{2}$ as the Jordan–Hölder constituents of a principal series representation $\pi$. Indeed, by [27, §1], $\pi_{n}$ is the Langlands quotient of the (unitarily normalized) parabolic induction of the character
(1) $\mu(\bar{\alpha},\beta,\alpha^{-1})=\lambda(\bar{\alpha})\nu(\beta)|\alpha|_{E}^{1/2},$
with $\pi_{2}$ the unique non-zero irreducible sub-representation. The sub and quotient are switched when $\mu$ is replaced by $\mu^{\mathsf{w}}(\bar{\alpha},\beta,\alpha^{-1})=\lambda(\bar{\alpha})\nu(\beta)|\alpha|_{E}^{-1/2}$, where $\mathsf{w}=\left(\begin{smallmatrix}&&1\\ &1&\\ 1&&\end{smallmatrix}\right)$ is the non-trivial element of the Weyl group of $G$. The Jacquet functor $\pi\to\pi_{N}$ is exact and it sends $\pi_{2}$ (resp. $\pi_{n}$) to $\mu\delta^{1/2}$ (resp. $\mu^{\mathsf{w}}\delta^{1/2}$), where $\delta(\bar{\alpha},\beta,\alpha^{-1})=|\alpha|_{E}^{2}$ is the modulus character. The extension
(2) $0\to\pi_{2}\to\pi=\operatorname{Ind}_{B}^{G}(\mu)\xrightarrow{\mathrm{pr}}\pi_{n}\to 0$
does not split since, by Frobenuis reciprocity, one has
$\operatorname{End}_{G}\left(\operatorname{Ind}_{B}^{G}(\mu)\right)=\operatorname{Hom}_{B}\left(\operatorname{Ind}_{B}^{G}(\mu),\mu\delta^{1/2}\right)=\operatorname{Hom}_{T}\left(\operatorname{Ind}_{B}^{G}(\mu)_{N},\mu\delta^{1/2}\right)\simeq\mathbf{C}.$
One knows by Rodier [25, Thm. 2] that the image of $\operatorname{Ind}_{B}^{G}(\mu)$ by a twisted (by a non-degenerate character of $N$) Jacquet functor, singling out generic representations, is a line. Since $\pi_{n}$ is non-generic (see [26, p.174]), the exactness of this functor implies that $\pi_{2}$ is generic.
As $\pi_{n}$ is non-tempered, the subspace $\pi_{2}$ consists of $f\in\pi$ such that for all $f^{\vee}\in\pi^{\vee}$ the matrix coefficient $g\mapsto\langle g\cdot f,f^{\vee}\rangle$ belongs to $\operatorname{L}^{2}(G)$. Conversely the following lemma holds.
Lemma 1.1.
Let $f\in\pi$. If $g\mapsto\langle g\cdot f,f^{\vee}\rangle$ belongs to $\operatorname{L}^{2}(G)$ for some $0\neq f^{\vee}\in\pi^{\vee}$, then $f\in\pi_{2}$.
Proof.
The dual of (2) is given by
$0\to\pi_{n}^{\vee}\to\pi^{\vee}=\operatorname{Ind}_{B}^{G}(\mu^{-1})\to\pi_{2}^{\vee}\to 0,$
and the irreducibility of $\pi_{2}$ and $\pi_{n}$ implies that $\pi_{n}^{\vee}=\{f^{\vee}\in\pi^{\vee}|\langle\pi_{2},f^{\vee}\rangle=0\}$. As $f^{\vee}\neq 0$, its $G$-span contains $\pi_{n}^{\vee}$, implying that the matrix coefficient $g\mapsto\langle g\cdot f,f^{\vee}\rangle$ belongs to $\operatorname{L}^{2}(G)$ for all $f^{\vee}\in\pi_{n}^{\vee}$. One deduces that
$g\mapsto\langle g\cdot f,f^{\vee}\rangle=\langle\mathrm{pr}(g\cdot f),f^{\vee}\rangle=\langle g\cdot\mathrm{pr}(f),f^{\vee}\rangle\in\operatorname{L}^{2}(G)$
As the irreducible $\pi_{n}$ is not a discrete series representation, this implies $\mathrm{pr}(f)=0$, i.e. $f\in\pi_{2}$. ∎
We will be mostly interested in the following $A$-packets having trivial central characters:
(3) $\Pi(\lambda)=\Pi(\lambda,\lambda_{|E^{1}}^{-1}).$
1.3. The Gross subgroup $K^{\prime\prime}$
In this subsection, $E$ is assumed ramified (hence $p$ is odd), $\mathcal{O}/\mathcal{P}=\mathbf{F}_{p}$ and $\mathcal{P}=(\xi)$. As $|\operatorname{PGL}(2,\mathbf{F}_{p})|=|\operatorname{SL}(2,\mathbf{F}_{p})|$ all vertices in the tree of $G$ have valence $p^{3}+1$. The reductive quotient $\overline{G}{}^{\circ}$ of $K^{\circ}$ is isomorphic to the orthogonal group $\mathrm{O}(3,\mathbf{F}_{p})$ with respect to the quadratic form represented by $\left(\begin{smallmatrix}&&1\\ &1&\\ 1&&\end{smallmatrix}\right)$. The adjoint action on matrices $\left(\begin{smallmatrix}y&x\\ z&-y\end{smallmatrix}\right)$ preserving the determinant $(y^{2}+xz)$ allows us to identify $\operatorname{PGL}(2,\mathbf{F}_{p})$ and $\mathrm{SO}(3,\mathbf{F}_{p})$ as follows:
(4) $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\mapsto\frac{1}{ad-bc}\begin{pmatrix}a^{2}&-ab&-b^{2}/2\\ -2ac&ad+bc&bd\\ -2c^{2}&2cd&d^{2}\end{pmatrix}$.
One has $\mathrm{O}(3,\mathbf{F}_{p})=\pm\mathbf{1}_{3}\cdot\mathrm{SO}(3,\mathbf{F}_{p})$. By the above description, $\mathrm{SO}(3,\mathbf{F}_{p})$ is generated by the set
$\left\{\left(\begin{smallmatrix}&&-1/2\\ &-1&\\ -2&&\end{smallmatrix}\right),\left(\begin{smallmatrix}a&&\\ &1&\\ &&a^{-1}\end{smallmatrix}\right),\left(\begin{smallmatrix}1&-b&-b^{2}/2\\ &1&b\\ &&1\end{smallmatrix}\right)\Big|a\in\mathbf{F}_{p}^{\times},b\in\mathbf{F}_{p}\right\}.$
Definition 1.2.
Let $K^{\prime\prime}$ be the index $2$ subgroup of $K^{\circ}$ defined as the inverse image of the subgroup of $\mathrm{O}(3,\mathbf{F}_{p})$ generated by $\mathbf{1}_{3}$ and the image of $\operatorname{PSL}(2,\mathbf{F}_{p})$. Let $I^{\prime\prime}=K^{\prime\prime}\cap K^{\prime}\subset I$.
As $E/\mathbf{Q}_{p}$ is a ramified quadratic extension with $p$ odd, a conjugate-symplectic character $\lambda$ of $E^{\times}$ is necessarily ramified and its restriction to $\mathbf{Z}_{p}^{\times}$ is given by its unique quadratic character. If $\lambda$ is tamely ramified, then its restriction to $\mathcal{O}^{\times}$ is also given by its unique quadratic character, and the equation $\lambda(\xi)^{2}=\lambda(-\xi\bar{\xi})=\lambda(-1)=(-1)^{(p-1)/2}$ shows that there are precisely two such characters.
Interested in determining a level for an element of the $A$-packet $\Pi(\lambda)$ considered in (3), we are indebted to B. Gross for generously sharing a suggestion that led to the following proposition.
Proposition 1.3.
Let $\lambda$ be a tamely ramified conjugate-symplectic character of $E^{\times}$ and let $\pi_{n}$ be the non-tempered member of the $A$-packet $\Pi(\lambda)$. Then $\dim\pi_{n}^{K^{\prime\prime}}=\dim\pi_{2}^{I^{\prime\prime}}=1$.
Proof.
As $\mu_{|(T\cap I^{\prime\prime})}=\mu^{\mathsf{w}}_{|(T\cap I^{\prime\prime})}=\mathbf{1}$, applying the Jacquet functor to the exact sequence of admissible $G$-representations (2) allows one to see that both $\pi_{2}$ and $\pi_{n}$ have non-trivial $I^{\prime\prime}$-invariants. Moreover, as $\big|B\backslash G/I^{\prime\prime}\big|=\big|(B\cap K^{\prime\prime})\backslash K^{\prime\prime}/I^{\prime\prime}\big|=2$, both $\pi_{2}^{I^{\prime\prime}}$ and $\pi_{n}^{I^{\prime\prime}}$ must be $1$-dimensional. By Iwasawa decomposition, the restriction of $\operatorname{Ind}_{B}^{G}(\mu)$ to $K^{\prime\prime}$ is given by $\operatorname{Ind}_{B\cap K^{\prime\prime}}^{K^{\prime\prime}}(\mu)$, hence the line $\operatorname{Ind}_{B}^{G}(\mu)^{K^{\prime\prime}}$ admits a basis $f$ uniquely characterized by $f_{|K^{\prime\prime}}=\mathbf{1}_{K^{\prime\prime}}$. It follows that $\dim\pi_{n}^{K^{\prime\prime}}+\dim\pi_{2}^{K^{\prime\prime}}=1$ and we will show that $\pi_{2}^{K^{\prime\prime}}=\{0\}$.
The line $(\operatorname{Ind}_{B}^{G}(\mu^{-1}))^{K^{\prime\prime}}$ admits a basis $f^{\vee}$ uniquely characterized by
$\langle v,f^{\vee}\rangle=\frac{1}{\sqrt{\operatorname{vol}(K^{\prime\prime})}}\int\limits_{K^{\prime\prime}}v(k)dk.$
[…]
Recall $\gamma=\left(\begin{smallmatrix}-\xi^{-1}&&\\ &1&\\ &&\xi\end{smallmatrix}\right)$ and let $\eta=\left(\begin{smallmatrix}\bar{u}&&\\ &1&\\ &&u^{-1}\end{smallmatrix}\right)$ where $u\in\mathcal{O}^{\times}$ is a fixed non-square element.
As $K^{\circ}=K^{\prime\prime}\coprod\eta K^{\prime\prime}$, Cartan decomposition for the special maximal compact $K^{\circ}$ yields:
$G=\coprod_{n\geqslant 0}\left(K^{\prime\prime}\gamma^{n}K^{\prime\prime}\right)\amalg\left(K^{\prime\prime}\gamma^{n}\eta K^{\prime\prime}\right).$
Since $\eta\cdot f=-f$ one deduces that $\langle\gamma^{n}\eta\cdot f,f^{\vee}\rangle=-\langle\gamma^{n}\cdot f,f^{\vee}\rangle$ and checking that $f\notin\pi_{2}$ amounts to proving the divergence of the numerical sequence with general term
$\operatorname{vol}(K^{\prime\prime}\gamma^{n}K^{\prime\prime})\big|\langle\gamma^{n}\cdot f,f^{\vee}\rangle\big|^{2}=[K^{\prime\prime}:(K^{\prime\prime}\cap\gamma^{n}K^{\prime\prime}\gamma^{-n})]\Big|\int\limits_{K^{\prime\prime}}f(k\gamma^{n})dk\Big|^{2}.$
By Iwahori decomposition one has $I^{\prime\prime}:(K^{\prime\prime}\cap\gamma^{n}K^{\prime\prime}\gamma^{-n})]=p^{2n-1}$ for all $n\geqslant 1$. As $\big|(\mu\delta^{1/2})(\gamma)\big|=p^{3/2}$ we are led to establish the divergence of the sequence with general term
$\Phi_{n}=p^{5n/2}\cdot\Big|\int\limits_{K^{\prime\prime}}f(\gamma^{-n}k\gamma^{n})dk\Big|.$
In view of the inequality $p>\sqrt{p}+1$ for $p\geqslant 3$, this will follow from the next lemma. ∎
Lemma 1.4.
For all $n\geqslant 1$ one has $\left|\int\limits_{K^{\prime\prime}{\smallsetminus}K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\right|\leqslant\operatorname{vol}(I^{\prime\prime})\cdot(\sqrt{p}+1)p^{-2n}$ and $\left|\int\limits_{K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\right|=\operatorname{vol}(I^{\prime\prime})\cdot p^{1-2n}$.
Proof.
The last row of an element $k\in K^{\prime\prime}$ is given by $(0,0,1)\cdot k=(c_{1}(k),c_{2}(k),c_{3}(k))\in\mathcal{O}^{3}{\smallsetminus}\mathcal{P}^{3}$. For $j\geqslant 0$ we let
$K^{\prime\prime}_{j}=\left\{k\in K^{\prime\prime}\big|c_{1}(k)\in\mathcal{P}^{j}\right\}\text{ and }K^{\prime\prime\times}_{j}=K^{\prime\prime}_{j}{\smallsetminus}K^{\prime\prime}_{j+1}.$
Note that $K^{\prime\prime}_{0}=K^{\prime\prime}$ and $K^{\prime\prime}_{1}=I^{\prime\prime}$. We use the partition $K^{\prime\prime}{\smallsetminus}K^{\prime\prime}_{2n}=K_{0}^{\prime\prime\times}\coprod K^{\prime\prime\times}_{1}\coprod\dots\coprod K^{\prime\prime\times}_{2n-1}$ to compute the first integral and $K^{\prime\prime}_{2n}=K^{\prime\prime\times}_{2n}\coprod K^{\prime\prime}_{2n+1}$ for the second.
For $j\geqslant 1$, using Iwahori decomposition, one finds that $I^{\prime\prime}:K^{\prime\prime}_{j}]=[I^{\prime\prime}\cap N^{-}:K^{\prime\prime}_{j}\cap N^{-}]=p^{j-1}$.
Given any $k\in K^{\prime\prime\times}_{j}(0\leqslant j\leqslant 2n)$, using the Iwasawa decomposition $G=\gamma^{\mathbf{Z}}NK^{\circ}=\gamma^{\mathbf{Z}}NK^{\prime}$, one finds that $\gamma^{n-j}k\gamma^{n}\in NK^{\circ}$, hence $|f(\gamma^{-n}k\gamma^{n})|\leqslant|\mu(\gamma^{j-2n})|=p^{\frac{3}{2}j-3n}$. Therefore
$\left|\int\limits_{K^{\prime\prime}{\smallsetminus}K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\right|\leqslant\operatorname{vol}(K^{\prime\prime}{\smallsetminus}I^{\prime\prime})\cdot p^{-3n}+\operatorname{vol}(I^{\prime\prime})\sum_{j=1}^{2n-1}p^{\frac{1}{2}j-3n}(p-1)=$
$\operatorname{vol}(I^{\prime\prime})\left(p^{\frac{1}{2}-2n}+p^{-2n}-p^{1-3n}-p^{\frac{1}{2}-3n}+([K^{\prime\prime}:I^{\prime\prime}]-1)p^{-3n}\right),$
proving the desired inequality, as $K^{\prime\prime}:I^{\prime\prime}]=p+1$ (obtained by going to the reductive quotient).
Since $\operatorname{vol}(I^{\prime\prime})\cdot p^{1-2n}=\operatorname{vol}(K^{\prime\prime}_{2n})$, in order to complete the proof of the lemma, it suffices to show that $f(\gamma^{-n}k\gamma^{n})$ is constant on $k\in K^{\prime\prime}_{2n}$. This is evident for $k\in K^{\prime\prime}_{2n+1}$, as then $k$ and $\gamma^{-n}k\gamma^{n}$ both belong to $I^{\prime\prime}$ and share same determinant and lower right coefficient $c_{3}$, implying that $f(\gamma^{-n}k\gamma^{n})=f(k)$. Miraculously, as one can see from (4), this remains true for $k\in K^{\prime\prime}_{2n}{\smallsetminus}K^{\prime\prime}_{2n+1}$ as well, i.e. even though $\gamma^{-n}k\gamma^{n}\in K^{\circ}{\smallsetminus}I$, the fact that $c_{3}(\gamma^{-n}k\gamma^{n})=c_{3}(k)$ still implies that $\gamma^{-n}k\gamma^{n}\in K^{\prime\prime}$. ∎
1.4. Higher Iwahori invariants via matrix coefficients
In this subsection, we assume that $E/\mathbf{Q}_{p}$ is unramified and we let $\lambda_{0}$ denote the unique quadratic unramified character of $E^{\times}$. By Iwasawa decomposition, the corresponding $\pi=\operatorname{Ind}_{B}^{G}(\mu_{0})$ has one dimensional invariants by any given maximal open compact subgroup $K$ of $G$. The following proposition states, depending on $K$, whether the $K$-invariant line belongs its sub-representation $\pi_{2}$ or maps non-trivially to the quotient $\pi_{n}$ (see §1.2 for notation). We recall that $K^{\circ}$ and $K^{\prime}$ are the two standard maximal compact subgroups, $K^{\circ}$ being the hyperspecial one, and that the standard Iwahori subgroup $I$ equals $K^{\circ}\cap K^{\prime}$.
Proposition 1.5.
One has $\dim\pi_{2}^{K^{\prime}}=\dim\pi_{n}^{K^{\circ}}=1$.
Proof.
Applying the Jacquet functor to the exact sequence (2) allows one to see that both $\pi_{2}$ and $\pi_{n}$ have $1$-dimensional II-invariants. As by Cartan decomposition $G$ is generated by $K$ and $\gamma$, hence by $K^{\circ}$ and $K^{\prime}$, it follows that exactly one amongst $\pi_{2}^{I}$ and $\pi_{n}^{I}$ is fixed by $K^{\circ}$, while the other one is fixed by $K^{\prime}$. By Iwasawa decomposition, the restriction to $K$ yields an isomorphism $\pi\xrightarrow{\sim}\operatorname{Ind}_{B\cap K}^{K}(\mu_{0})$, under which a basis $f_{K}$ of $\pi^{K}$ is mapped to $\mathbf{1}_{K}$. Moreover, the line $(\pi^{\vee})^{K}$ admits a basis $f_{K}^{\vee}$ characterized by
$\langle f,f_{K}^{\vee}\rangle=\frac{1}{\sqrt{\operatorname{vol}(K)}}\int\limits_{K}f(k)dk.$
[…]
$\sqrt{\operatorname{vol}(K\gamma^{n}K)}\langle\gamma^{n}\cdot f_{K},f_{K}^{\vee}\rangle=\sqrt{[K:(K\cap\gamma^{-n}K\gamma^{n})]}\int\limits_{K}f_{K}(k\gamma^{n})dk.$
Using Iwahori decomposition for all $n\geqslant 1$ we have $K:(K\cap\gamma^{-n}K\gamma^{n})]/[K:I]=p^{4n-3}$ (resp. $p^{4n-1}$), where $K=K^{\circ}$ (resp. $K^{\prime}$). Since $(\mu_{0}\delta^{1/2})(\gamma)=-p^{3}$ we have $f_{K}\in\pi_{2}$ if, and only if,
(5) $(\Phi^{K}_{n})_{n}\in\operatorname{L}^{2}(\mathbf{Z}_{\geqslant 0})\text{, where }\Phi^{K}_{n}=p^{5n}\cdot\int\limits_{K}f_{K}(\gamma^{-n}k\gamma^{n})dk.$
The proof of the Proposition is then completed by the following Lemma. ∎
Lemma 1.6.
The quantity $p^{n}\cdot\Phi^{K^{\prime}}_{n}$ is independent of $n\geqslant 1$, in particular $(\Phi^{K^{\prime}}_{n})_{n}\in\operatorname{L}^{2}(\mathbf{Z}_{\geqslant 0})$.
Proof.
The last row of an element $k\in K^{\prime}$ is given by $(0,0,1)\cdot k=(p\cdot c_{1}(k),p\cdot c_{2}(k),c_{3}(k))$ with $c_{2}(k)\in\mathcal{O}$ and $(c_{1}(k),c_{3}(k))\in(\mathcal{O}\times\mathcal{O}){\smallsetminus}(\mathcal{P}\times\mathcal{P})$. For $j\geqslant 0$ we let
$K^{\prime}_{j}=\left\{k\in K^{\prime}\Big|c_{1}(k)\in\mathcal{P}^{j}\right\}\text{ and }K^{\prime\times}_{j}=K^{\prime}_{j}{\smallsetminus}K^{\prime}_{j+1}.$
To compute the above integral we use the partition $K^{\prime}=K_{0}^{\prime\times}\coprod K^{\prime\times}_{1}\coprod\dots\coprod K^{\prime\times}_{2n-1}\coprod K^{\prime}_{2n}$.
First, we compute the volume of $K^{\prime}_{j}$, for $j\geqslant 1$. Using Iwahori decomposition one finds that
$K^{\prime}:K^{\prime}_{j}]=\frac{[K^{\prime}:I]}{[K^{\prime}_{j}:I\cap K^{\prime}_{j}]}[I:I\cap K^{\prime}_{j}]=\frac{[K^{\prime}:I]}{[K^{\prime}\cap N:I\cap N]}[I\cap N^{-}:K^{\prime}_{j}\cap N^{-}]=c_{0}\cdot p^{j+2\left[\frac{j}{2}\right]}.$
Next we observe that by Iwasawa decomposition, for all $k\in K^{\prime}_{2n}$ one has $c_{2}(k)\in\mathcal{P}^{n}$, i.e. $\gamma^{-n}k\gamma^{n}\in N\cdot K^{\prime}$, and therefore $f_{K}(\gamma^{-n}k\gamma^{n})=1$.
Using again Iwasawa decomposition, one checks that for $0\leqslant j\leqslant 2n-1$ and for every $k\in K^{\prime\times}_{j}$ one has $p^{n-j}c_{2}(k)\in\mathcal{O}^{\times}$, hence $\gamma^{-n}k\gamma^{n}\in\gamma^{j-2n}N\cdot K^{\prime}$ and $f_{K}(\gamma^{-n}k\gamma^{n})=(-p^{3})^{j-2n}$.
Therefore $\frac{1}{\operatorname{vol}(K^{\prime})}\int\limits_{K^{\prime}}f_{K}(\gamma^{-n}k\gamma^{n})dk=\frac{1}{[K^{\prime}:K^{\prime}_{2n}]}+p^{-6n}\sum_{j=0}^{2n-1}(-1)^{j}p^{3j}\frac{1}{[K^{\prime}:K^{\prime\times}_{j}]}=$
$c_{0}\cdot p^{-6n}\left(p^{2n}+c_{0}^{-1}-\sum_{i=1}^{n}p^{6i-3}(p^{-4i+3}-p^{-4i})+\sum_{i=0}^{n-1}p^{6i}(p^{-4i}-p^{-4i-1})\right)=p^{-6n}(1+c_{0}).\qed$
Remark 1.7.
Alternatively, one could use MacDonald’s formula for zonal spherical functions to see that $\pi_{2}$ does not admit non-zero vectors fixed by the hyperspecial maximal open compact subgroup $K^{\circ}$. Indeed, [16, §5.5] allows one to write $\Phi^{K^{\circ}}_{n}$ from (5), up to a non-zero constant, as
$\Gamma_{\mu_{0}}\cdot\mu_{0}(\gamma^{-n})+\Gamma_{\mu_{0}^{\mathsf{w}}}\cdot\mu_{0}^{\mathsf{w}}(\gamma^{-n}),\text{ with }\Gamma_{\nu}=\frac{1-p^{2}\cdot\nu(\gamma^{-2})}{1-\nu(\gamma^{-2})}\cdot\frac{1-p^{2}\cdot\nu(\gamma^{-1})}{1-\nu(\gamma^{-1})},$
where the two factors in $\Gamma_{\nu}$ correspond respectively to the positive roots $\zeta$ and $2\zeta$ of $G$ (see §1.1). As $\mu_{0}(\gamma^{-1})=\mu_{0}^{\mathsf{w}}(\gamma)=-p^{-1}$, one has $\Gamma_{\mu_{0}}=0\neq\Gamma_{\mu_{0}^{\mathsf{w}}}$, hence the sequence $(\Phi^{K^{\circ}}_{n})_{n\geqslant 0}$ is not $\operatorname{L}^{2}$. This also shows, in passing, that $\pi_{n}$ is not square integrable.
As a consequence we obtain the following lower bound, in the unramified case.
Corollary 1.8.
For $r\geqslant 1$ and for $\pi_{n}\in\Pi(\lambda_{0})=\Pi(\lambda_{0},\mathbf{1})$, one has $\dim\left(\pi_{n}^{I_{2r}}\right)\geqslant r+1$.
Proof.
Let $f$ be a basis of $\pi_{n}^{K^{\circ}}$ (see Proposition 1.5). For all $r\in\mathbf{Z}$, $\pi_{n}$ contains a (unique) line fixed by $\gamma^{r}K^{\circ}\gamma^{-r}$ (having $\gamma^{r}\cdot f_{K^{\circ}}$ as basis) and moreover, the stabilizer in $G$ of that line is $\gamma^{r}K^{\circ}\gamma^{-r}$. We claim that the vectors $f,\gamma\cdot f,\dots,\gamma^{r}\cdot f\in\pi_{n}$ are linearly independent. Indeed, if $f$ was a linear combination of the remaining $r$ vectors then it would be fixed by $\bigcap_{1\leqslant j\leqslant r}\gamma^{i}K^{\circ}\gamma^{-i}$ which is not contained in $K^{\circ}$. As any of these $(r+1)$ vectors is fixed by $I_{2r}$, the claim follows.
A similar claim holds for $\pi_{2}$ and can be proven exactly as above, using $K^{\prime}$ instead of $K^{\circ}$. ∎
While the unramified $A$-packet $\Pi(\lambda_{0})$ would suffice for our global applications when $D\equiv 3\pmod{8}$, the case of discriminants $D\equiv 7\pmod{8}$ would require the use of certain tamely ramified $A$-packets $\Pi(\lambda)$ and providing explicit levels for them is the object of the next subsection.
1.5. Intertwinings and exponential sums
Our arithmetic applications will require to show existence of non-zero $I_{r}$-invariants, for some $r\in\mathbf{Z}_{\geqslant 1}$, in certain ramified $A$-packets. This is delicate because of the lack of a new-vector theory for non-generic representations $\pi$ (see Remark 1.11). Also Casselman’s result asserting that $\pi^{I_{r}}$ surjects onto $\pi_{N}^{T\cap I_{r}}$ is inconclusive here as the latter vanishes, in contrast with the situation in [11] where the open compact is a pro-$p$-Iwahori subgroup of a sufficiently deep level. We will instead employ explicit methods and prove in Theorem 1.9 that $\pi_{n}$ has non-zero invariants by $K_{T}$, which contains a conjugate of $I_{3}$. It is relatively straightforward to determine all such vectors in the full induced representation, but it becomes a thorny comptutation to find a non-square integrable matrix coefficient as in §1.4. By another result of Casselman, matrix coefficients can be expressed in terms of the corresponding ones in the Jacquet module, given here by an explicit scalar product on $\pi_{N}\simeq\mathbf{C}^{2}$. Making this actually work requires non-vanishing of the second coordinate of the image of a $K_{T}$-invariant vector under the Jacquet functor which, once verified, leads directly to the result we seek.
We reduce the computation of the Jacquet functor to a precise statement about the intertwining operator at the level of finite reductive groups. The latter involves showing non-vanishing of some explicit exponential sums, bringing out the arithmetic nature of the problem. Although these sums seem extremely hard to compute individually, we manage to deduce it by precisely computing a suitable average of such sums corresponding to the trace of the finite intertwining operator.
Theorem 1.9.
Assume that $p$ odd and $E/\mathbf{Q}_{p}$ unramified. Fix a character $\chi:E^{1}\twoheadrightarrow\mathbf{F}^{1}_{p^{2}}\to\mathbf{C}^{\times}$ such that $\chi^{3}\neq\mathbf{1}$ and let $\lambda=\lambda_{0}\cdot\chi_{E}$. Letting $\pi_{n}$ denote the non-tempered representation of the Arthur packet $\Pi(\lambda)$, one has $\dim\pi_{n}^{K_{T}}=1$.
Lemma 1.10.
Consider the character $\mu_{\lambda}$ from (1) attached to $\lambda$ and trivial on $E^{1}$. One has
$B\backslash G/K_{T}=\overline{B}\backslash\overline{G}/\overline{T}=\left\{\mathbf{1}_{3},\mathsf{w},\sigma_{\infty}=\left(\begin{smallmatrix}1&0&0\\ 0&1&0\\ \xi&0&1\end{smallmatrix}\right),\sigma_{y}=\left(\begin{smallmatrix}1&0&0\\ -1&1&0\\ \xi y-\frac{1}{2}&1&1\end{smallmatrix}\right)\Big|y\in\mathbf{F}_{p}\right\}.$
For any non-trivial $\chi$, the $K_{T}$-invariant functions in $\operatorname{Ind}_{B}^{G}(\mu_{\lambda})$ are supported by the double cosets of $\{\sigma_{y}|y\in\mathbf{F}_{p}\}$ and, in addition when $\chi$ is quadratic, by the double coset of $\sigma_{\infty}$.
Proof.
By the Iwasawa decomposition, restriction to $K$ yields an isomorphism $\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\xrightarrow{\sim}\operatorname{Ind}_{B\cap K}^{K}(\mu_{\lambda})$. As $\mu_{\lambda|B\cap K}$ factors through a character $\overline{\mu}_{\lambda}$ of $\overline{B}$, the space of $K_{1}$-invariants in $\operatorname{Ind}_{B\cap K}^{K}(\mu_{\lambda})$ is naturally identified with $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})$, and therefore its subspace of $K_{T}$-invariants is given by $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})^{\overline{T}}$. By definition, a double coset $\overline{B}\sigma\overline{T}$ supports a non-zero function contributing to the latter space if, and only if, $\overline{\mu}_{\lambda}$ is trivial on $\overline{B}\cap\sigma\overline{T}\sigma^{-1}$ (this is in accordance with Mackey’s Theorem for finite groups). As $\overline{\mu}_{\lambda}$ is ramified this is never the case for $\sigma=\mathbf{1}_{3}$ nor for $\sigma=\mathsf{w}$, while a matrix computation shows that $\sigma=\sigma_{\infty}$ works if, and only if, $\chi$ is quadratic. For $y\in\mathbf{F}_{p}$ and $k\in K_{T}$, writting $(k\bmod{p})=\operatorname{diag}\left(\bar{\alpha},\beta,\alpha^{-1}\right)$, we find
$(\sigma_{y}k\sigma_{y}^{-1}\bmod{p})=\left(\begin{smallmatrix}\bar{\alpha}&0&0\\ \beta-\bar{\alpha}&\beta&0\\ *&\beta-\alpha^{-1}&\alpha^{-1}\end{smallmatrix}\right)$
ensuring the triviality of $\mu_{\lambda}$ on $B\cap\sigma_{y}K_{T}\sigma_{y}^{-1}$, as in forces $\bar{\alpha}\equiv\beta\pmod{p}$ on such elements. ∎
Recall that $\pi_{n}$ is the Langlands quotient of $\operatorname{Ind}_{B}^{G}(\mu_{\lambda})$ whose other Jordan-Hölder constituent is the discrete series $\pi_{2}$. As taking invariants by an open compact subgroup is an exact functor in the category of admissible representations, Lemma 1.10 implies that
(6) $\dim\pi_{n}^{K_{T}}+\dim\pi_{2}^{K_{T}}=\begin{cases} p&\text{ if } \chi^{2}\neq\mathbf{1},\\ p+1&\text{if } \chi\text{ quadratic}.\end{cases}$
Remark 1.11.
To show $\pi_{n}^{K_{T}}\neq\{0\}$ one could instead try computing $\dim\pi_{2}^{K_{T}}$. Miyauchi’s theory [20, 21] of conductors with respect to the paramodular groups $K^{\prime}_{r}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&p^{-r}\mathcal{O}\\ p^{r}\mathcal{O}&\mathcal{O}^{\times}&\mathcal{O}\\ p^{r}\mathcal{O}&p^{r}\mathcal{O}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$ asserts that the level of the generic $\pi_{2}$ is given by its conductor, while $\pi_{c}$ and the ramified $\pi_{n}$ have no level, i.e. they have no non-zero invariants by $K^{\prime}_{r}$ for any $r$. When $\chi$ is the quadratic character, the $L$-parameter $(\lambda\otimes\mathrm{St})\oplus\mathbf{1}$ has conductor $2$, therefore the generic member $\pi_{2}$ in this $L$-packet has $1$-dimensional invariants by $K^{\prime}_{2}$, hence also by $\gamma^{-1}K^{\prime}_{2}\gamma=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{P}&\mathcal{O}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{P}\\ \mathcal{O}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G\supset K_{T}$. For other ramified $\chi$’s, $\pi_{2}$ has an invariant line by $K^{\prime}_{3}$, having the same volume as $K_{T}$ but not conjugate to it, thus not settling the non-vanishing of $\pi_{2}^{K_{T}}$, let alone computing its dimension.
In order to show that $\pi_{n}$ admits non-zero $K_{T}$-invariants, it would be enough to consider the image of a $K_{T}$-invariant vector from Lemma 1.10 under the Jacquet functor and show that its second coordinate does not vanish. As well known, this functor is given by
$\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\longrightarrow\mathbf{C}\cdot\mu_{\lambda}\delta^{1/2}\oplus\mathbf{C}\cdot\mu_{\lambda}^{\mathsf{w}}\delta^{1/2},\quad f\mapsto(f(1),(Mf)(1))$
where the standard intertwining operator $M:\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\to\operatorname{Ind}_{B}^{G}(\mu_{\lambda}^{\mathsf{w}})$ is defined via analytic continuation, as follows. For $s\in\mathbf{C}$, letting $\mu_{\lambda,s}=\mu_{\lambda}\cdot\delta^{s/2}$, the intertwining operator
$M_{s}:\operatorname{Ind}_{B}^{G}(\mu_{\lambda,s})\to\operatorname{Ind}_{B}^{G}(\mu_{\lambda,s}^{\mathsf{w}}),\quad f_{s}\mapsto\int\limits_{N}f_{s}(\mathsf{w}n\cdot)dn$
is absolutely convergent for $\Re(s)\gg 0$ and $G$-equivariant. Moreover, for any section $f_{s}\in\operatorname{Ind}_{B}^{G}(\mu_{\lambda,s})$ such that for all $g\in G$ the function $f_{s}(g)$ is analytic in $s\in\mathbf{C}$, the function $(M_{s}(f_{s}))(g)$, a priori only defined for $\Re(s)\gg 0$, is a rational function in $p^{-s}$, hence extends to a meromorphic function on all of $\mathbf{C}$ with only possibly a finite number of poles independent of $f$ and of $g$. In fact, it continues as an intertwining operator, i.e.
$(M_{s}(f_{s}))(gg^{\prime})=(M_{s}(f_{s}(\cdot g^{\prime}))(g),\text{ for all }g,g^{\prime}\in G.$
We refer to [1, §1] for more detail and proofs. We will computing explicitly the right hand side of
$M(f_{0})=\lim_{s\to 0}M_{s}(f_{s})$
as a rational function in $p^{-s}$, simultaneously justifying the absence of pole at $s=0$.
For any $y\in\mathbf{F}_{p}$, Lemma 1.10 yields a unique function $f_{y}\in\left(\operatorname{Ind}_{B}^{G}(\mu_{\lambda})\right)^{K_{T}}$ supported on $B\sigma_{y}K_{T}$ and normalized by $f_{y}(\sigma_{y})=1$. In addition, when $\chi$ is quadratic, we let $f_{\infty}$ be the unique function supported on $B\sigma_{\infty}K_{T}$ and normalized by $f_{\infty}(\sigma_{\infty})=1$. Consider a $K^{\circ}$-flat section $f_{y,s}$ passing through $f_{y,0}=f_{y}$.
Proposition 1.12.
For all $y,y^{\prime}\in\mathbf{F}_{p}\cup\{\infty\}$ and $\Re(s)\gg 0$, we have
$M_{s}(f_{y^{\prime},s})(\sigma_{y})=\chi(-1)\frac{(1-p)p^{-3-2s}}{1-p^{-1-2s}}\delta_{y,y^{\prime}}+p^{-3}\sum_{[z,x]\in N(\mathbf{F}_{p})}f_{y^{\prime},s}\left(\mathsf{w}[z,x]\sigma_{y}\right).$
Proof.
Recall that $z,x]=\left(\begin{smallmatrix}1&-\bar{z}&u\\ &1&z\\ &&1\end{smallmatrix}\right)$, where $u=\xi\cdot x-\tfrac{z\bar{z}}{2}$, so that $u+\bar{u}+z\bar{z}=0$. We first compute the intertwining integral over
[…]
a key observation being that $z,x]\in N(\mathbf{Z}_{p})$ if and only if $u\in\mathcal{O}$. It follows that for $z,x]\in N(\mathbf{Q}_{p}){\smallsetminus}N(\mathbf{Z}_{p})$ both $1/u$ and $\bar{z}/u$ belong to $\mathcal{P}$, yielding an explicit Iwasawa decomposition:
$\mathsf{w}n=\mathsf{w}[z,x]=\begin{pmatrix}&&1\\ &1&z\\  1&-\bar{z} &u\end{pmatrix}=\begin{pmatrix}\bar{u}^{-1}&\bar{z}/u&1\\ &-\bar{u}/u&z\\  &&u\end{pmatrix}\begin{pmatrix}1&&\\ z/\bar{u}&1&\\  1/u&-\bar{z}/u &1\end{pmatrix}$.
As the last term belongs to the principal congruence subgroup $K_{1}$ which is normalized by $\sigma_{y}\in K^{\circ}$ and contained in $K_{T}$, we deduce that $f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})=\begin{cases}\mu_{\lambda,s}(\bar{u}^{-1},-\bar{u}/u,u)&\text{ if } y^{\prime}=y,\\ 0&\text{ if } y^{\prime}\neq y.\end{cases}$ Hence for $u\in p^{-m}\mathcal{O}^{\times}$, we have $f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})=\mu_{\lambda,s}(\gamma)^{-m}\chi^{3}_{E}(\bar{u}^{-1})=(-1)^{m}\cdot p^{-(3+2s)m}\chi^{3}_{E}(u)$.
For $m$ odd, noticing further that $\chi_{E}(u)=\chi_{E}(\xi)=\chi(-1)$, we find
$\int\limits_{[p^{(1-m)/2}\mathcal{O},p^{-m}\mathbf{Z}_{p}^{\times}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=-\chi(-1)\cdot p^{-(3+2s)m}p^{m-1}(p^{m}-p^{m-1}),\text{ hence}$
(7) $\sum\limits_{m\geqslant 1\text{ odd}} \int\limits_{[p^{(1-m)/2}\mathcal{O},p^{-m}\mathbf{Z}_{p}^{\times}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=\frac{-\chi(-1)(p^{-2-2s}-p^{-3-2s})}{1-p^{-2-4s}}.$
We similarly find
(8) $\sum\limits_{m\geqslant 2\text{ even}}\int\limits_{[p^{(2-m)/2}\mathcal{O},p^{-m}\mathbf{Z}_{p}^{\times}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=\frac{\chi(-1)(p^{-4-4s}-p^{-5-4s})}{1-p^{-2-4s}}.$
Finally, for $m\geqslant 2$ even we compute
$\int\limits_{[p^{-\frac{m}{2}}\mathcal{O}^{\times},p^{-m}\mathbf{Z}_{p}]}f_{y,s}(\mathsf{w}n\sigma_{y})dn=p^{-(3+2s)m}\cdot p^{m-2}p^{m-1} \sum\limits_{[z_{0},x_{0}]\in\mathbf{F}_{p^{2}}^{\times}\times\mathbf{F}_{p}}\chi^{3}_{E}(z_{0}\bar{z}_{0}-2\xi x_{0}).$
As $\chi^{3}\neq\mathbf{1}$, $\sum\limits_{(z_{0},x_{0})\in\mathbf{F}_{p^{2}}^{\times}\times\mathbf{F}_{p}}\chi^{3}\left(\frac{z_{0}\bar{z}_{0}-2\xi x_{0}}{z_{0}\bar{z}_{0}+2\xi x_{0}}\right)=(p^{2}-1)\sum\limits_{x_{0}\in\mathbf{F}_{p}}\chi^{3}\left(\frac{1-\xi x_{0}}{1+\xi x_{0}}\right)=-\chi(-1)(p^{2}-1)$ yields
(9) $\sum\limits_{m\geqslant 2\text{ even}}\int\limits_{[p^{-\frac{m}{2}}\mathcal{O}^{\times},p^{-m}\mathbf{Z}_{p}]}f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})dn=\frac{-\chi(-1)(p^{-3-4s}-p^{-5-4s})}{1-p^{-2-4s}}.$
Putting (7), (8) and (9) together, we find $\int\limits_{N(\mathbf{Q}_{p}){\smallsetminus}N(\mathbf{Z}_{p})}f_{y,s}(\mathsf{w}n\sigma_{y})dn=-\chi(-1)\frac{(p^{-2-2s}-p^{-3-2s})}{1-p^{-1-2s}}$.
Finally, as $f_{y^{\prime},s}$ is invariant by $K_{T}\supset K_{1}\supset N(p\mathbf{Z}_{p})$, we have
$\int\limits_{N(\mathbf{Z}_{p})}f_{y^{\prime},s}(\mathsf{w}n\sigma_{y})dn=p^{-3}\sum_{[z,x]\in N(\mathbf{Z}_{p})/N(p\mathbf{Z}_{p})}f_{y^{\prime},s}\left(\mathsf{w}[z,x]\sigma_{y}\right).\qed$
The coefficients of the intertwining automorphism of $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})^{\overline{T}}$ in the basis $(f_{y})_{y}$ are given by
$M_{y,y^{\prime}}=\sum_{[z,x]\in\mathbf{F}_{p^{2}}\times\mathbf{F}_{p}}f_{y^{\prime},0}\left(\mathsf{w}[z,x]\sigma_{y}\right).$
Corollary 1.13.
Letting $\mathscr{M}_{0}$ (resp. $\mathscr{M}_{0}^{\mathsf{w}}$) denote the matrices of $M_{0}$ (resp. $M_{0}^{\mathsf{w}}$) in the bases $(f_{y})_{y}$ and $(f^{\mathsf{w}}_{y})_{y}$, we have
$\mathscr{M}:=(M_{y,y^{\prime}})_{(y,y^{\prime})}=\chi(-1)p\cdot I_{p}+p^{3}\cdot\mathscr{M}_{0}=-\chi(-1)p^{2}\cdot I_{p}+p^{3}\cdot\mathscr{M}_{0}^{\mathsf{w}}.$
Proof.
The first equality is obtained by letting $s$ go to 0 in Proposition 1.12, while the second one is obtained by applying the same process to $\mu_{\lambda}^{\mathsf{w}}$, noting that $\mu_{\lambda,s}^{\mathsf{w}}=\mu_{\lambda,s-1}$, in particular $\overline{\mu}_{\lambda}=\overline{\mu}_{\lambda}^{\mathsf{w}}$. ∎
Lemma 1.14.
For all $y\in\mathbf{F}_{p}$, one has $\sum\limits_{y^{\prime}\in\mathbf{F}_{p}}M_{y,y^{\prime}}=\sum\limits_{y\in\mathbf{F}_{p}}M_{y,y^{\prime}}=\chi(-1)p$. More importantly
(10) $\operatorname{Tr}(\mathscr{M})=\begin{cases}-\chi(-1)p&\text{ if } \chi^{2}\neq\mathbf{1}\quad(\text{and} \chi^{3}\neq\mathbf{1}),\\ 0&\text{ if } \chi\text{ quadratic}.\end{cases}$
Proof.
As $f_{y^{\prime}}(\mathsf{w}\sigma_{y})=0$, we can assume that $u\neq 0$. Letting $v=1/u$ and $w=1-\bar{z}/u$, we have
$\mathsf{w}[z,x]\sigma_{y}=\begin{pmatrix}-\tfrac{1}{2}+\xi y&1&1\\ -1-\tfrac{z}{2}+\xi yz&1+z&z\\  1+\bar{z}-\tfrac{u}{2}+\xi yu&u-\bar{z}&u\end{pmatrix}=\begin{pmatrix}\bar{u}^{-1}&*&*\\ &-\bar{u}/u&*\\  &&u\end{pmatrix}\begin{pmatrix}1&&\\ -\bar{w}&1&\\  v-w+\tfrac{1}{2}+\xi y&w&1\end{pmatrix}$
Hence $f_{y^{\prime}}(\mathsf{w}[z,x]\sigma_{y})\neq 0$ if and only if $v-w+\tfrac{1}{2}+\xi y=w\bar{w}(-\tfrac{1}{2}+\xi y^{\prime})\neq 0$, i.e.,
[…]
in which case it equals $\chi_{E}(u^{3}w)=\chi_{E}(w)\cdot\chi^{3}_{E}(\bar{v})$. Noting $(w-1)(\bar{w}-1)+v+\bar{v}=0$, so that for a fixed $w\neq 1$, the ratio $\tfrac{\bar{v}}{v}=\tfrac{u}{\bar{u}}=\frac{\xi x-(w-1)(\bar{w}-1)/2}{-\xi x-(w-1)(\bar{w}-1)/2}$ describes $\mathbf{F}^{1}_{p^{2}}{\smallsetminus}\{-1\}$, we find
$\sum_{y\in\mathbf{F}_{p}}M_{y,y^{\prime}}=\sum_{y^{\prime}\in\mathbf{F}_{p}}M_{y,y^{\prime}}=$
$=\sum_{w\neq 1} \chi_{E}(w)\sum_{x\in\mathbf{F}_{p}} \chi^{3}_{E}\left(\xi x-(w-1)(\bar{w}-1)/2\right)+\sum_{x\in\mathbf{F}_{p}^{\times}} \chi^{3}(-1)$
$=\sum_{w\neq 1} \chi_{E}(w)(-\chi(-1))+(p-1)\chi(-1)=p\chi(-1).$
To compute $\operatorname{Tr}(\mathscr{M})$ we observe that when $w\notin\mathbf{F}^{1}_{p^{2}}$, (11) imposes no condition on $v$, as we are summing over all $y\in\mathbf{F}_{p}$, and that when $w\in\mathbf{F}^{1}_{p^{2}}$, it imposes no condition on $y$ while determines uniquely $v=\tfrac{w-\bar{w}}{2}-\tfrac{(w-1)(\bar{w}-1)}{2}=w-1$ . As $v\neq 0$ implies $w\neq 1$, we find
$\sum_{y\in\mathbf{F}_{p}}M_{y,y}=-\chi(-1)\sum_{w\in\mathbf{F}_{p^{2}}^{\times}{\smallsetminus}\mathbf{F}^{1}_{p^{2}}} \chi_{E}(w)+p\sum_{w\in\mathbf{F}^{1}_{p^{2}}{\smallsetminus}\{1\}} \chi_{E}(w(\bar{w}-1)^{3})$
$=\chi(-1)\sum_{w\in\mathbf{F}^{1}_{p^{2}}} \chi^{2}(w)+p\sum_{w\in\mathbf{F}^{1}_{p^{2}}{\smallsetminus}\{1\}}\chi(-\bar{w})=\begin{cases}-p\chi(-1)&\text{ if } \chi^{2}\neq\mathbf{1},\\ \chi(-1)&\text{ if } \chi\text{ quadratic}.\end{cases}$
Finally, for $\chi$ quadratic, letting $w=-\tfrac{\bar{z}}{u}$, we find $\mathsf{w}[z,x]\sigma_{\infty}=\left(\begin{smallmatrix}\bar{u}^{-1}&*&*\\ &-\bar{u}/u&*\\  &&u\end{smallmatrix}\right)\left(\begin{smallmatrix}1&&\\ -\bar{w}&1&\\  u^{-1}+\xi&w&1\end{smallmatrix}\right)$. This implies that $f_{\infty}(\mathsf{w}[z,x]\sigma_{\infty})=0$ unless $w=z=0$ in which case it equals $\chi_{E}(\xi y)$, with $y\bar{y}=1+\xi^{-2}x^{-1}\neq 0$. Then one finds $M_{\infty,\infty}=\sum\limits_{x\in\mathbf{F}_{p}^{\times}}f_{\infty}\left(\mathsf{w}[0,x]\sigma_{\infty}\right)=-\chi(-1)$. Combining with the second case of the above equation, one finds $\operatorname{Tr}(\mathscr{M})=M_{\infty,\infty}+\sum\limits_{y\in\mathbf{F}_{p}}M_{y,y}=0$. ∎
Proof of Theorem 1.9.
As $\overline{\mu}_{\lambda}=\overline{\mu}_{\lambda}^{\mathsf{w}}$, Frobenius reciprocity yields $\operatorname{End}_{\overline{G}}\left(\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})\right)=\mathbf{C}^{2}$. Hence $\operatorname{Ind}_{\overline{B}}^{\overline{G}}(\overline{\mu}_{\lambda})$ is a direct sum of two irreducible distinct $\overline{G}$-representations, denoted $\overline{\pi}_{n}$ and $\overline{\pi}_{2}$, on each of which the intertwining endomorphism $\mathscr{M}$ acts scalarly by Schur’s Lemma.
Furthermore, as $\overline{\pi}_{n}^{\overline{T}}$ and $\overline{\pi}_{2}^{\overline{T}}$ are naturally identified with $\pi_{n}^{K_{T}}=\ker(\mathscr{M}_{0}^{\mathsf{w}})$ and $\pi_{2}^{K_{T}}=\ker(\mathscr{M}_{0})$ (see (2)), it follows from Corollary 1.13 that the above mentioned eigenvalues are respectively $\chi(-1)p^{2}$ and $\chi(-1)p\cdot I_{p}$. Formula (10) can thus be rewritten as
$p\cdot\dim\pi_{n}^{K_{T}}-\dim\pi_{2}^{K_{T}}=-\frac{\chi(-1)}{p}\operatorname{Tr}(\mathscr{M})=\begin{cases}1&\text{ if } \chi^{2}\neq\mathbf{1}\quad(\text{and} \chi^{3}\neq\mathbf{1}),\\ 0&\text{ if } \chi\text{ quadratic},\end{cases}$
which combined with (6) yields $\dim\pi_{n}^{K_{T}}=1$. ∎
2. Galois representations for $3$-folds of Picard type
From this point onwards, we will use global notations from the introduction. The local results of the previous section can be applied to the completion $E$ of $M$ at any prime number which does not split in that field. We denote by $\mathbf{A}_{f}$ the ring of finite adeles of $\mathbf{Q}$, so that $\mathbf{A}=\mathbf{A}_{f}\times\mathbf{R}$.
2.1. Abelian $3$-folds and Shimura surfaces of Picard type
Let $k$ be any field containing $M$. Consider an abelian $3$-fold $A/k$ together with an injection $\iota^{0}:M\hookrightarrow\mathrm{End}^{0}(A/k)=\mathrm{End}(A/k)\otimes\mathbf{Q}$, or equivalently with an injection $\iota$ of an order of $M$ into $\mathrm{End}(A/k)$, the most important for us case being when $\iota^{0}$ comes from $\iota:\mathcal{O}_{M}\hookrightarrow\mathrm{End}(A/k)$.
A polarization on $A$ is an isogeny $\theta:A\to A^{\vee}$, where $A^{\vee}$ denotes the dual abelian variety. By positivity, since $k$ is a field, the Rosati involution induced by $\theta$ on $\iota^{0}(M)$ is given by the complex conjugation (see [22, §21]). A polarization is called principal, if it is an isomorphism, and it can always be acquired over a finite extension of $k$.
By an abelian $3$-fold of Picard type over $k$ we will mean a principally polarized abelian variety over $k$ of dimension $3$ having multiplication by $\mathcal{O}_{M}$ defined over $k$.
The action of $M$ splits the $3$-dimensional $k$-vector space $\operatorname{Lie}(A/k)$ in a direct sum of two sub-spaces: one on which the actions of $M$ and $k$ agree, and one on which they differ by the complex conjugation, the pair of their dimension being called the signature.
[…]
which is free of rank $3$ over $\mathbf{A}_{M,f}$. Given a polarization $\theta:A\to A^{\vee}$, the Weil pairing endows $V_{f}A$ with a non-degenerate skew-hermitian form, i.e., a non-degenerate alternating pairing
$\langle\cdot,\cdot\rangle_{A}:V_{f}A\times V_{f}A\to\mathbf{A}_{f}$
such that $\langle a\cdot v,v^{\prime}\rangle_{A}=\langle v,\bar{a}\cdot v^{\prime}\rangle_{A}$ for all $a\in M$. If $\theta$ is principlal, then $\langle\cdot,\cdot\rangle_{A}$ is a perfect pairing.
Let $(V,\langle\cdot,\cdot\rangle)$ be a $3$-dimensional (non-degenerate) hermitian space over $M$. The unitary similitude group $\widetilde{G}=\mathrm{GU}(V)$ is the reductive group over $\mathbf{Q}$ characterized by the property that
$\widetilde{G}(R)=\{g\in\operatorname{GL}(V\otimes_{\mathbf{Q}}R)\mid\forall v,v^{\prime}\in V\otimes_{\mathbf{Q}}R,\langle g(v),g(v^{\prime})\rangle=\nu(g)\langle v,v^{\prime}\rangle\},$
for any $\mathbf{Q}$-algebra $R$, where $\nu:\mathrm{GU}(V)\to\mathbf{G}_{m,\mathbf{Q}}$ is a homomorphism whose kernel is the unitary group $G=\operatorname{U}(V)$. Since any hermitian form in $3$ variables over a non-archimedean local field is isotropic, the group $\widetilde{G}(\mathbf{Q}_{p})$ is unique up to isomorphism, while at infinity $\langle\cdot,\cdot\rangle$ is uniquely determined by its signature, hence there are two possibilities for $\widetilde{G}(\mathbf{R})$ (as opposite signatures define isomorphic groups). Hasse’s Principle applied to the semi-simple simply connected derived group $G^{1}=\mathrm{SU}(V)$ implies then that, up to an isomorphism, there exists a unique quasi-split unitary group, denoted $\widetilde{G}=\mathrm{GU}(2,1)$, and a unique definite unitary group, denoted $\mathrm{GU}(3)$.
We will now introduce the Shimura variety for $\widetilde{G}$ represented by the matrix $\left(\begin{smallmatrix}&&1\\ &\sqrt{-D}&\\ -1&&\end{smallmatrix}\right)$. The homomorphism
$\tilde{h}:\mathrm{Res}^{\mathbf{C}}_{\mathbf{R}}\mathbf{G}_{m,\mathbf{R}}\to\widetilde{G}_{\mathbf{R}}\text{ , }z\mapsto\begin{pmatrix}\Re(z)&0&\Im(z)\\ 0&z&0\\ -\Im(z)&0&\Re(z)\end{pmatrix}$
of $\mathbf{R}$-algebraic groups satisfies Deligne’s axioms [8] for a Shimura variety, hence for any open compact subgroups $\widetilde{K}$ of $\widetilde{G}(\mathbf{A}_{f})$ one can consider the Shimura surface
[…]
where $\mathcal{H}$ is identified with the $\widetilde{G}(\mathbf{R})$-conjugacy classes of $\tilde{h}$. By a fundamental result of Shimura, $Y_{\widetilde{K}}$ admits a canonical model over the reflex field $M$. The connected components of $Y_{\widetilde{K}}$ are Picard modular surfaces.
For the anisotropic form $\mathrm{GU}(3)$, one can analogously define Shimura sets which are finite and therefore will not alter the uniformity of our results in §3.
[…]
Let $(A,\iota,\theta)$ be an abelian $3$-fold of Picard type over a number field $k$. The action of $\operatorname{Gal}_{k}$ on the adelic Tate module $V_{f}A$, together with the choice of $\eta$ as above, yields a continuous homomorphism:
(12) $\rho_{A,f}:\operatorname{Gal}_{k}\longrightarrow\widetilde{G}(\mathbf{A}_{f}).$
[…]
In analogy with the index $2$ Gross subgroups $\widetilde{K}^{\prime\prime}_{p}$ of maximal compacts $K^{\circ}_{p}$ at ramified primes introduced in §1.3, we consider the open compact subgroup
(13) $\widetilde{K}^{\prime\prime}=\widetilde{K}^{\prime\prime}_{D}\prod_{p\nmid D}\widetilde{K}_{p}^{\circ}\subset\widetilde{G}(\mathbf{A}_{f}),$
where $\widetilde{K}^{\prime\prime}_{D}$ is defined as the kernel of the composed homomorphism
(14) $\prod_{p\mid D}\widetilde{K}_{p}^{\circ}\twoheadrightarrow\prod_{p\mid D}\widetilde{K}_{p}^{\circ}/\widetilde{K}^{\prime\prime}_{p}=\prod_{p\mid D}\{\pm 1\}\xrightarrow{\Pi}\{\pm 1\}.$
Let $(A,\iota,\theta)$ be a principally polarized abelian $3$-fold of Picard type over $k$. For $v$ the prime of $M$ above $p\mid D$, the action of $\operatorname{Gal}_{k}$ on $A[v]$ factors through $\mathrm{GO}(3,\mathbf{F}_{p})$. Using the exceptional isomorphism $\mathrm{PGO}(3,\mathbf{F}_{p})\xrightarrow{\sim}\mathrm{SO}(3,\mathbf{F}_{p})\xrightarrow{\sim}\operatorname{PGL}(2,\mathbf{F}_{p})$, one defines its projectivization
(15) $\widetilde{\rho}_{A,p}:\operatorname{Gal}_{k}\to\operatorname{PGL}(2,\mathbf{F}_{p}).$
Taking quotient by the unique index two subgroup $\operatorname{PSL}(2,\mathbf{F}_{p})$ of $\operatorname{PGL}(2,\mathbf{F}_{p})$ yields a canonical homomorphism $\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$ and we let $\varepsilon_{A,D}=\prod_{p\mid D}\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$.
Using the observation made after (12), the points in $Y_{\widetilde{K}^{\prime\prime}}(k)$ corresponds precisely to an abelian $3$-fold AA over $k$ of Picard type having trivial $\varepsilon_{A,D}$.
2.2. Complex reflexions and families of abelian threefolds of Picard type
It is important to observe that even though for $\widetilde{K}\subset\widetilde{K}^{\circ}$ each point of $Y_{\widetilde{K}}(\mathbf{C})$ is associated to an abelian $3$-fold of Picard type, there does not exist such a family over the entire $Y_{\widetilde{K}}(\mathbf{C})$ unless there are no points with extra automorphisms, in which case $Y_{\widetilde{K}}(\mathbf{C})$ would be a fine moduli space. In our cases of interest $\widetilde{K}$ is not neat, and therefore $Y_{\widetilde{K}}$ is not a fine moduli space. In this subsection we prove that there exists a natural family of abelian $3$-folds of Picard type over $Y_{\widetilde{K}^{\prime\prime}}$ minus a finite number of points (see (13)). This will be crucially used in the proof of Theorem B.
Recall that an element $\gamma$ of the discrete subgroup $\Gamma=\widetilde{G}(\mathbf{Q})\cap\widetilde{K}\widetilde{G}(\mathbf{R})$ has a fixed point in $\mathcal{H}$ if, and only if, $\gamma$ has finite order (this is because the stabilizers in $\widetilde{G}(\mathbf{R})$ of points in $\mathcal{H}$ are maximal compact subgroups). Such a $\gamma$ is called elliptic if it only fixes an isolated point in $\mathcal{H}$, otherwise it is called a complex reflexion. By [17, Cor. 4.5.10], the set of singular points of $Y_{\Gamma}$ is finite and consists of isolated fixed points of elliptic elements of $\Gamma$, all defined over a finite extension of $M$. Moreover, smooth points at which universal cover $\mathcal{H}\to Y_{\Gamma}$ is not étale are fixed points of complex reflexions, i.e.  order $2$ elements in $\Gamma$ fixing hyperbolic planes (see [17, §4.5]).
[…]
Lemma 2.1.
The group $\Gamma^{\prime\prime}$ does not contain any complex reflexions.
Proof.
Consider a complex reflexion $\gamma\in\Gamma^{\prime\prime}$ as an endomorphism of the Hermitian space $M^{3}$ having signature $(2,1)$. As the eigenspaces of $\gamma$ are mutually orthogonal, at most one such eigenspace could contain a negative line (corresponding to a point in $\mathcal{H}$). This if $\gamma$ fixes more that one point of $\mathcal{H}$, it necessarily fixes a hyperbolic line in $\mathcal{H}$. The corresponding endomorphism of $M^{3}$ has a mutually orthogonal eigenplane and eigenline (both $M$-rational), forcing the eigenvalues to be in $\mathcal{O}_{M}^{\times}=\{\pm 1\}$ (as $D>4$) and not all equal. It follows, that for any $p\mid D$, the image of $g_{f}^{-1}\gamma g_{f}\in\widetilde{K}^{\circ}$ into the projectivization of the reductive quotient of $\widetilde{K}_{p}^{\circ}$ belongs to the image under the adjoint isomorphism $\operatorname{PGL}(2,\mathbf{F}_{p})\xrightarrow{\sim}\mathrm{PGO}(3,\mathbf{F}_{p})$ of an element represented by a matrix having both eigenvalues $1$ and 1-1. In particular its image in $\operatorname{PGL}(2,\mathbf{F}_{p})/\operatorname{PSL}(2,\mathbf{F}_{p})=\{\pm 1\}$ equals $\left(\tfrac{-1}{p}\right)$. As $\prod_{p\mid D}\left(\tfrac{-1}{p}\right)=\left(\tfrac{-1}{D}\right)=-1$ it follows from (13) that $g_{f}^{-1}\gamma g_{f}\notin\widetilde{K}^{\prime\prime}$, i.e., $\gamma\notin\Gamma^{\prime\prime}$. ∎
Proposition 2.2.
There exists a finite extension $k$ of $M$ and a natural family of abelian $3$-folds of Picard type over $Y_{\widetilde{K}^{\prime\prime}}\times_{M}k$ minus a finite number of $k$-rational elliptic points.
Proof.
We claim that there is a family of abelian $3$-folds of Picard type over any open subset $U$ of $Y_{\widetilde{K}^{\prime\prime}}$ which contains no point with a non-trivial stabilizer. By [18, §2.3.4], the moduli stack $S_{\widetilde{K}^{\prime\prime}}$ associated to this problem is an algebraic stack (for the étale topology), locally of finite type over the base which we may take to $\operatorname{Spec}(M)$. Moreover, by [18, §A.7.5], there is a canonical surjective morphism $\phi$ from $S_{\widetilde{K}^{\prime\prime}}$ to the associated coarse moduli space $S_{\widetilde{K}^{\prime\prime}}$], which in our notations is $Y_{\widetilde{K}^{\prime\prime}}$. By [18, §7], $S_{\widetilde{K}^{\prime\prime}}$] is an algebraic space and even a quasi-projective scheme. Moreover, by a general property of moduli stacks (see [23, Chap. 7]), $\phi$ is an isomorphism over the locus $U$ where there is no non-trivial automorphism, by which we mean it has no infinitesimal automorphism; analytically, this corresponds to points of $Y_{\widetilde{K}}(\mathbf{C})$ having no non-trivial stabilizers. Now $U$ is a priori an open subscheme of $S_{\widetilde{K}^{\prime\prime}}$], but since it is where $\phi$ is an isomorphism, we get a canonical open $U\hookrightarrow S_{\widetilde{K}^{\prime\prime}}$. This map tautologically yields the desired family $f:A\to U$ of abelian varieties of Picard type. In view of Lemma 2.1 and the discussion preceding it, after possibly enlarging $k$, we may consider $U$ be the complementary open in $Y^{\prime\prime}$ of its finitely many elliptic points. ∎
2.3. Étale fundamental groups and Mumford–Tate groups
Let $k$ be a number field containing $M$ over which the connected component of $Y_{\widetilde{K}}$ are defined. Fix a connected component $Y$ of $Y_{\widetilde{K}}\times_{M}k$ and a smooth open $U$ of $Y$ endowed with an abelian scheme $f:A\to U$ of Picard type. Denote by $\eta$ the generic point of the smooth surface $U$. Fixing a closed geometric point $\bar{x}$ of $U$ the étale fundamental group sits in the middle of a short exact sequence
(16) $1\to\Pi_{1}(U_{\bar{k}},\bar{x})\to\Pi_{1}(U,\bar{x})\to\operatorname{Gal}_{k}\to 1.$
The morphism $f:A\to U$ being proper and smooth, one can consider the étale sheaf $\mathrm{R}^{1}f_{*}\mathbf{Z}_{\ell}$ on $U$. As $U$ is geometrically connected, we have $\Pi_{1}(U,\bar{x})\simeq\Pi_{1}(U,\bar{\eta})$ and the latter acts on
$(\mathrm{R}^{1}f_{*}\mathbf{Z}_{\ell})_{\bar{\eta}}=\operatorname{H}^{1}(A_{\bar{\eta}},\mathbf{Z}_{\ell})=(T_{\ell}A_{\eta})^{\vee},$
yielding a continuous representation
$\operatorname{Gal}(\bar{\eta}/\eta)\twoheadrightarrow\Pi_{1}(U,\bar{x})\xrightarrow{\rho_{U,\ell}}\operatorname{Aut}_{\mathbf{Z}_{\ell}}(T_{\ell}A_{\eta}).$
Any closed point $x\in U(k)$ yields a section $s_{x}:\operatorname{Gal}_{k}\to\Pi_{1}(U,\bar{x})$ of (16) allowing one to consider
$\rho_{x,\ell}=\rho_{U,\ell}\circ s_{x}:\operatorname{Gal}_{k}\to\operatorname{Aut}_{\mathbf{Z}_{\ell}}(T_{\ell}A_{\eta}).$
Finally for any closed curve $C\subset U$ defined over $k$, there is a natural map $\Pi_{1}(C,\bar{x})\to\Pi_{1}(U,\bar{x})$ whose composition with $\rho_{U,\ell}$ is denoted $\rho_{C,\ell}$. As $f:A\to U$ is of Picard type, for any $x\in C(k)$,
$\Gamma_{x}=\mathrm{im}(\rho_{x,\ell})\subset\Gamma_{C}=\mathrm{im}(\rho_{C,\ell})\subset\Gamma_{U}=\mathrm{im}(\rho_{U,\ell})\subset K^{\circ}_{\ell}.$
By a series of results of Cadoret–Tamagawa (see [3, 4]), the set $C_{\rho}$ of all $x\in C(k)$ for which $\Gamma_{x}$ is not open in $\Gamma_{C}$ is finite and the index $\Gamma_{C}:\Gamma_{x}]$ is uniformly bounded for $x\in C(k){\smallsetminus}C_{\rho}$.
Recall that the Mumford–Tate group $\operatorname{MT}(A)$ of a polarized abelian variety $A$ over $\mathbf{C}$ is the smallest connected reductive subgroup of $\operatorname{GL}(\operatorname{H}_{1}(A,\mathbf{Q}))$ over $\mathbf{Q}$, whose $\mathbf{R}$-points contain the image of the associated $\mathbf{R}$-morphism $h:\mathbf{C}^{\times}\to\operatorname{GL}(\operatorname{H}_{1}(A(\mathbf{C}),\mathbf{R}))$ coming from the Hodge decomposition. If $A$ is defined over a field $k$ finitely generated over $\mathbf{Q}$, then the image $\Gamma_{\ell}$ of $\operatorname{Gal}_{k}$ acting on $T_{\ell}A$ is an $\ell$-adic Lie group. Denoting $\mathfrak{g}_{\mathbf{Z}_{\ell}}$ its Lie algebra, it is a theorem of Deligne [9, Chap. I.2] that $\mathfrak{g}_{\mathbf{Q}_{\ell}}=\mathfrak{g}_{\mathbf{Z}_{\ell}}\otimes_{\mathbf{Z}_{\ell}}\mathbf{Q}_{\ell}\subset\operatorname{Lie}(\operatorname{MT}(A)\otimes_{\mathbf{Q}}\mathbf{Q}_{\ell})$ and the Mumford–Tate conjecture, known for abelian varieties of dimension at most $3$, asserts that they are equal (see e.g. [5]).
As the Mumford–Tate group of the (generic point of the) universal family $f:A\to U$ is given by $\widetilde{G}=\mathrm{GU}(2,1)$, it follows from Deligne [9, Cor. 6.2] that the Mumford–Tate group of any abelian $3$-fold of Picard type is a reductive subgroup of $\widetilde{G}$. We have the following trichotomy.
Lemma 2.3.
Let $\mathfrak{g}$ be a reductive Lie subalgebra of $\mathfrak{gu}(3,\mathbf{Q}_{\ell})$. If $\mathfrak{g}^{\prime}\subset\mathfrak{su}(3,\mathbf{Q}_{\ell})$ denotes the semi-simple part of $\mathfrak{g}$, exactly one of the following holds:
  1. (i)
    $\mathfrak{g}^{\prime}=\{0\}$, i.e., $\mathfrak{g}$ is abelian,
  2. (ii)
    $\mathfrak{g}^{\prime}$ is a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$,
  3. (iii)
    $\mathfrak{g}^{\prime}=\mathfrak{su}(3,\mathbf{Q}_{\ell})$.
Proof.
If $\mathfrak{g}^{\prime}_{\overline{\mathbf{Q}}_{\ell}}=\{0\}$, then $\mathfrak{g}^{\prime}=\{0\}$, whereas if $\mathfrak{g}^{\prime}_{\overline{\mathbf{Q}}_{\ell}}=\mathfrak{sl}(3,\overline{\mathbf{Q}}_{\ell})$, then $\mathfrak{g}^{\prime}=\mathfrak{su}(3,\mathbf{Q}_{\ell})$ for dimension reasons. In the remaining cases, using the well known fact that any proper non-zero semi-simple Lie subalgebra of $\mathfrak{sl}(3,\overline{\mathbf{Q}}_{\ell})$ is isomorphic to $\mathfrak{sl}(2,\overline{\mathbf{Q}}_{\ell})$, we deduce that $\mathfrak{g}^{\prime}$ is a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$. ∎
If $A$ is of CM type (resp. admits a CM factor), then $\operatorname{Lie}(\operatorname{MT}(A)\otimes_{\mathbf{Q}}\mathbf{Q}_{\ell})$ is of the first (resp. second) type in Lemma 2.3. Also clearly $\rho_{A,\ell}$ is potentially abelian (resp. is potentially reducible) if, and only if, $\mathfrak{g}_{\mathbf{Q}_{\ell}}$ is of the first (resp. second) type in Lemma 2.3. The following proposition bridges the two sides.
Proposition 2.4.
Let $A$ be an abelian $3$-fold of Picard type over a number field $k$. Then $\rho_{A,\ell}$ is potentially abelian (resp. is potentially reducible) if, and only if, $A$ is of CM type (resp. admits a CM factor).
Proof.
One implication is clear. For the other, after extending scalars to $\overline{\mathbf{Q}}_{\ell}$ and after possibly replacing $k$ by a finite extension, we may assume that $\rho_{A,\ell}$ is potentially reducible, i.e. it contains (as a direct factor by Faltings) a character $\chi_{\ell}:k^{\times}\backslash\mathbf{A}_{k}^{\times}\to\overline{\mathbf{Q}}_{\ell}^{\times}$. Being a sub-representation of $\rho_{A,\ell}$, $\chi_{\ell}$ is unramified outside a finite set of places, and its restriction to decomposition groups at places above $\ell$ is Hodge-Tate with weights 0 and 1-1. In addition it is pure of weight 1-1. By Weil, $\chi_{\ell}$ corresponds to an algebraic Hecke character $\chi:k^{\times}\backslash\mathbf{A}_{k}^{\times}\to\mathbf{C}^{\times}$ whose infinity component is necessarily of the form $\mathrm{N}_{\Phi^{\prime}}\circ\mathrm{N}_{k/L^{\prime}}$ where $\mathrm{N}_{\Phi^{\prime}}$ is the partial norm given by a CM type $\Phi^{\prime}$ for a CM field $L^{\prime}\subset k$. By [31, Lem. 2] replacing $(L^{\prime},\Phi^{\prime})$ by its double reflex yields the same infinite component, hence we may and do assume that $(L^{\prime},\Phi^{\prime})$ is primitive, i.e., coincides with the reflex of a CM field $L$ endowed with a CM type $\Phi$. Further replacing $k$ by a finite (abelian) extension one can assume that $\chi_{f}$ takes values in $L^{\times}$. By Casselman (see [31, Thm. 6]), there exists an abelian variety $A^{\prime}$ defined over $k$ which is CM of type $(L,\Phi)$ and such that $\rho_{A^{\prime},\ell}=\chi_{\ell}$, hence
$\operatorname{Hom}_{\operatorname{Gal}_{k}}(\rho_{A,\ell},\rho_{A^{\prime},\ell})\neq\{0\}.$
By Faltings, one deduces that $\operatorname{Hom}_{k}(A,A^{\prime})\neq\{0\}$, hence $A$ contains a non-trivial CM quotient. If $A$ is of CM type we’re done. If not, then there exists an abelian variety $A^{\prime\prime}$ which is not of CM type and such that $A$ is isogenous to $A^{\prime}\times A^{\prime\prime}$, i.e., $V_{\ell}A=V_{\ell}A^{\prime}\oplus V_{\ell}A^{\prime\prime}$. Furthermore since $\operatorname{Hom}_{k}(A^{\prime},A^{\prime\prime})=\{0\}$, one can show that the isogeny is $\mathcal{O}_{M}$-linear. Hence $A^{\prime\prime}$ is also of Picard type, from which one deduces that we’re in the second case in Lemma 2.3, in particular $\rho_{A,\ell}$ is not potentially abelian. ∎
3. Uniform irreducibility of Galois images
3.1. A lemma on surfaces with isolated singularities
We denote by $q(X)$ the irregularity of a projective algebraic surface $X$, given by the dimension of $\operatorname{H}^{1}(X,\mathcal{O}_{X})$. If $X$ is smooth and projective over $\mathbf{C}$, then $q(X)=\dim\operatorname{H}^{0}(X,\Omega^{1}_{X})$.
Let $X$ be a projective irreducible algebraic surface over $\mathbf{C}$ with isolated singularities, i.e., such that there exists a smooth open $j:U\hookrightarrow X$ whose complement $Z=X{\smallsetminus}U$ consists of finitely many closed points. There exists a smooth resolution
$\phi:\widetilde{X}\to X$
such that $\phi^{-1}(Z)$ is a divisor with normal crossings with $\phi$ restricting to an isomorphism from $\phi^{-1}(U)$ onto $U$. Thus we get an injection $\widetilde{j}:U\hookrightarrow\widetilde{X}$ such that $j=\phi\circ\widetilde{j}$, and we denote by
$\widetilde{j}^{\ast}:\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})\to\operatorname{H}^{1}(U,\mathbf{Q})$
the pullback homomorphism on Betti cohomology. By [6, Thm. 3.2.5(iii)] we know that $\widetilde{j}^{\ast}$ is a homomorphism of mixed Hodge structure, with $\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})$ being pure of weight $1$.
Lemma 3.1.
The map $\widetilde{j}^{\ast}$ is an isomorphism, in particular $\operatorname{H}^{1}(U,\mathbf{Q})$ is a pure weight $1$ Hodge structure and $q(X)=\dim\operatorname{H}^{0}(U,\Omega^{1}_{U})$.
Proof.
Let $\mathrm{IH}^{\bullet}(X,\mathbf{Q})$ denote the middle intersection cohomology of $X$. Since
$\dim(X)-1>0=\dim(Z)$
by [10, Thm. 5.4.12] $j^{\ast}:\mathrm{IH}^{1}(X,\mathbf{Q})\to\mathrm{IH}^{1}(U,\mathbf{Q})$ is an isomorphism, while $\widetilde{j}^{\ast}$ is injective. Moreover by Corollary 5.4.11 and Proposition 5.4.4 in loc.cit. $\phi^{\ast}:\mathrm{IH}^{1}(X,\mathbf{Q})\to\mathrm{IH}^{1}(\widetilde{X},\mathbf{Q})=\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})$ is an embedding, while $\mathrm{IH}^{1}(U,\mathbf{Q})=\operatorname{H}^{1}(U,\mathbf{Q})$. This is summarized in the following commutative diagram:
${\mathrm{IH}^{1}(X,\mathbf{Q})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{\ast}}\scriptstyle{\sim}$$\scriptstyle{j^{\ast}}$${\mathrm{IH}^{1}(U,\mathbf{Q})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{j}^{\ast}}$${\operatorname{H}^{1}(U,\mathbf{Q})}$
It immediately follows that $\widetilde{j}^{\ast}$ is an isomorphism and $q(\widetilde{X})=\dim\operatorname{H}^{0}(\widetilde{X},\Omega^{1}_{\widetilde{X}})=\dim\operatorname{H}^{0}(U,\Omega^{1}_{U})$. Finally q(X)=q(X~)q(X)=q($\widetilde${X}) as the irregularity is a birational invariant. ∎
3.2. Irregularity of Picard modular surfaces
Let $z\mapsto\bar{z}$ be the non-trivial automorphism of $M/\mathbf{Q}$. Put $M^{1}=\{z\in M^{\times}\mid z\bar{z}=1\}$ viewed as an algebraic torus over $\mathbf{Q}$ and denote by $\mathbf{A}_{M}^{1}$ its adelic points.
[…]
$Y_{K}(\mathbf{C})=G(\mathbf{Q})\backslash\left(\mathcal{H}\times G(\mathbf{A}_{f})/K\right),$
[…]
Proposition 3.2.
The irregularity of any connected component of the minimal compactification $Y_{\widetilde{K}}^{\ast}$ of $Y_{\widetilde{K}}$ is given by the formula
(17) $q(Y_{\Gamma}^{\ast})=\sum_{(\lambda,\nu)\in\Xi/\widehat{\pi_{0}(Y_{K})}}\sum_{\pi_{f}\in\Pi_{f}(\lambda,\nu)}\dim(\pi_{f}^{K})\frac{1+W(\lambda\nu^{-1}_{M})(-1)^{s(\pi_{f})}}{2},\text{ where }$
  • $\Xi$ is the set of pairs $(\lambda,\nu)$ of a Hecke character $\lambda$ of $M$ whose restriction to $\mathbf{Q}$ is $\left(\tfrac{\cdot}{D}\right)$ and of a character $\nu$ of $\mathbf{A}_{M}^{1}/M^{1}$, such that
[…]
  • $\Pi_{f}(\lambda,\nu)$ is the finite part of the global Arthur packet from §1.2,
  • $W(\lambda\nu_{M}^{-1})\in\{\pm 1\}$ is the global root number, where $\nu_{M}(z)=\nu(z/\bar{z})$ for $z\in\mathbf{A}_{M}^{\times}$,
  • $s(\pi_{f})$ the number of finite places $v$ at which $\pi_{v}\simeq\pi_{c}(\lambda_{v},\nu_{v})$, and
  • $\mu\in\widehat{\pi_{0}(Y_{K})}$ acts freely on $\Xi$ by sending $(\lambda,\nu)$ to $(\lambda\mu_{M},\nu\mu)$.
Proof.
We first observe that the complement of $Y_{\Gamma}$ in $Y_{\Gamma}^{\ast}$ consists of finitely many points, the cusps. Furthermore, as explained in §2.2, $Y_{\Gamma}$ admits finitely many isolated singularities all of which are elliptic points. Thus there exists a smooth open $U_{K}$ of the normal projective surface $Y_{K}^{\ast}$ whose complement consists of finitely many closed points. Lemma 3.1 applied component-wise to $U_{K}$ yields $q(Y_{K}^{\ast})=\dim\operatorname{H}^{0}(U_{K},\Omega^{1}_{U_{K}})$. Let $K^{\prime}$ be any normal finite index torsion free subgroup of $K$, e.g. the intersection with the principal congruence subgroup of level $3$ (see [11, Lem. 1.4]). By Koecher’s Principle, as $Y_{K^{\prime}}{\smallsetminus}U_{K^{\prime}}$ has codimension at least $2$ in $Y_{K^{\prime}}$, we have
$\dim\operatorname{H}^{0}(U_{K},\Omega^{1}_{U_{K}})=\dim\operatorname{H}^{0}(U_{K^{\prime}},\Omega^{1}_{U_{K^{\prime}}})^{K/K^{\prime}}=\dim\operatorname{H}^{0}(Y_{K^{\prime}},\Omega^{1}_{Y_{K^{\prime}}})^{K/K^{\prime}},$
where $U_{K^{\prime}}$ is the inverse image of $U_{K}$ under the natural projection $Y_{K^{\prime}}\to Y_{K}$. Taking invariants by the finite group $K/K^{\prime}$ in Rogawski’s formula [11, (14)] for $q(Y_{K^{\prime}}^{\ast})=\dim\operatorname{H}^{0}(Y_{K^{\prime}},\Omega^{1}_{Y_{K^{\prime}}})$ yields
$\dim\operatorname{H}^{0}(U_{K},\Omega^{1}_{U_{K}})=\sum_{(\lambda,\nu)\in\Xi}\sum_{\pi_{f}\in\Pi(\lambda_{f},\nu_{f})}\dim(\pi_{f}^{K})\frac{1+W(\lambda\nu_{M}^{-1})(-1)^{s(\pi_{f})}}{2}$
(there is a misprint in loc. cit. where one should read $(1+W(\lambda\nu_{M}^{-1})(-1)^{s(\pi_{f})})$ instead of $(W(\lambda\nu_{M}^{-1})+(-1)^{s(\pi_{f})})$; it also uses the inverse notation for the base change $\nu_{M}$). The presence of this root number translates the fact that for $\pi_{f}\in\Pi(\lambda_{f},\nu_{f})$ and $\pi_{\infty}$ the unique non-tempered holomorphic representation in the local Arthur packet $\Pi(\lambda_{\infty},\nu_{\infty})$, $\pi=\pi_{f}\otimes\pi_{\infty}$ is automorphic if, and only if, $W(\lambda\nu_{M}^{-1})=(-1)^{s(\pi_{f})}$. Both $\dim(\pi_{f}^{K})$ and $1+W(\lambda\nu_{M}^{-1})(-1)^{s(\pi_{f})}$ being preserved by the action of $\widehat{\pi_{0}(Y_{K})}$, one deduces the desired formula for $q(Y_{\Gamma}^{\ast})$ as in [11, (15)]. ∎
3.3. Twists of canonical characters and root numbers
Hecke characters $(\lambda,\nu)\in\Xi$ whose local components at each finite place have ‘minimal’ ramification are intimately related to the canonical characters studied by Gross and Rohrlich. They play a pivotal role in our production of automorphic forms contributing to the irregularity of the Picard modular surfaces of low level. We will now briefly recall some of their properties under the running assumption that $D>3$ is odd. Consider the character $\lambda_{\infty}(z)=\bar{z}\cdot|z|^{-1}$ of $M_{\infty}^{\times}\simeq\mathbf{C}^{\times}$ and let $\lambda_{f}:\widehat{\mathcal{O}}_{M}^{\times}\to\mathbf{C}^{\times}$ be a continuous character whose restriction to $\mathcal{O}_{M,p}^{\times}$ is given by the unique quadratic character
$\mathcal{O}_{M,p}^{\times}\to\mathbf{F}_{p}^{\times}\xrightarrow{\left(\tfrac{\cdot}{p}\right)}\{\pm 1\},$
for all $p$ dividing $D$, and is trivial otherwise. As $\left(\tfrac{-1}{D}\right)=-1$, it follows that $\lambda_{\infty}$ and $\lambda_{f}$ agree on $\mathcal{O}_{M}^{\times}=\{\pm 1\}$. The finiteness of the ideal class group $\mathscr{C}\!\ell_{M}=\mathbf{A}_{M}^{\times}/M^{\times}\widehat{\mathcal{O}}_{M}^{\times}M_{\infty}^{\times}$ guarantees that the resulting character of $M^{\times}\widehat{\mathcal{O}}_{M}^{\times}M_{\infty}^{\times}$ extends to a character $\lambda$ of $\mathbf{A}_{M}^{\times}$ and clearly two such extensions must differ by an ideal class character. As by construction the restriction of $\lambda_{f}$ to $\widehat{\mathbf{Z}}^{\times}$ agrees with the quadratic Dirichlet character $\left(\tfrac{\cdot}{D}\right)$ viewed as a character of $\mathbf{A}^{\times}/\mathbf{Q}^{\times}\mathrm{Nm}(\mathbf{A}_{M}^{\times})=\operatorname{Gal}(M/\mathbf{Q})$ and $\mathbf{A}^{\times}=\mathbf{Q}^{\times}\widehat{\mathbf{Z}}^{\times}\mathbf{R}_{+}^{\times}$, it follows that the restriction of $\lambda$ to $\mathbf{A}^{\times}$ equals $\left(\tfrac{\cdot}{D}\right)$, i.e., $\lambda$ is conjugate-symplectic. Such a character is called canonical and we will denote it by $\lambda_{c}$, remembering that it is only unique up to a multiplication by a character of $\mathscr{C}\!\ell_{M}$. The root number $W(\lambda_{c}^{3})=-W(\lambda_{c})=\left(\tfrac{-2}{D}\right)$ is $1$ if, and only if, $D\equiv 3\pmod{8}$ (see [28]).
Assume henceforth that $\det(K)=\widehat{\mathcal{O}}_{M}^{1}$, so that $\pi_{0}(Y_{K})=\mathscr{C}\!\ell_{M}^{1}:=\mathbf{A}_{M}^{1}/M^{1}\widehat{\mathcal{O}}_{M}^{1}M_{\infty}^{1}$, and that $\widehat{\mathcal{O}}_{M}^{1}$ embeds centrally into $K$, the central character $\omega=\nu\cdot\lambda_{|M^{1}}$ of any $\pi$ contributing to (17) has to be everywhere unramified, i.e.,
(18) $q(Y_{\Gamma}^{\ast})=\frac{1}{|\mathscr{C}\!\ell_{M}^{1}|}\sum_{\chi\in\Xi^{1},\omega\in\widehat{\mathscr{C}\!\ell_{M}^{1}}}\sum_{\pi_{f}\in\Pi_{f}\left(\lambda_{c}\chi_{M},{\lambda^{-1}_{c}}_{\!\!\!\!\!|M^{1}}\chi^{-2}\omega\right)}\dim(\pi_{f}^{K})\frac{1+W(\lambda_{c}^{3}\chi^{3}_{M})(-1)^{s(\pi_{f})}}{2},$
where $\Xi^{1}$ denotes the set of finite order characters of $\mathbf{A}_{M}^{1}/M^{1}$ (see [28, (3)] for the fact that multiplication by an ideal class character does not change the root number).
If $3$ does not divide the class number $h:=|\mathscr{C}\!\ell_{M}|$, then the action of $\mu\in\widehat{\mathscr{C}\!\ell_{M}^{1}}$ sending $(\chi,\omega)$ to $(\chi\mu,\omega\mu^{3})$ allows one to twist out the central character $\omega$ and obtain the simpler formula
(19) $q(Y_{\Gamma}^{\ast})=\sum_{\chi\in\Xi^{1}}\sum_{\pi_{f}\in\Pi_{f}(\lambda_{c}\chi_{M})}\dim(\pi_{f}^{K})\frac{1+W(\lambda_{c}^{3}\chi^{3}_{M})(-1)^{s(\pi_{f})}}{2}.$
Successfully applying (18) requires one to understand how root numbers behave under twisting. As we are interested in creating irregularity at level $\Gamma^{\prime\prime}_{0}(\ell^{r})$, we focus on characters $\chi$ which are only ramified at the fixed inert prime $\ell$.
Lemma 3.3.
For $\chi\in\Xi^{1}$ of Artin conductor $\ell^{a(\chi)}$ and $\lambda$ an odd power of a canonical character,
$W(\lambda\chi_{M})=(-1)^{a(\chi)}\chi_{\ell}(-1)W(\lambda).$
Proof.
Using the factorization of root numbers $W=\prod\limits_{v}W_{v}$, it suffices to prove that
[…]
(21) $W_{v}(\lambda_{v}\chi_{M,v})$ $=W_{v}(\lambda_{v}),\,\text{ for all } v\neq\ell,$
where the local factors are defined using the standard additive character $\psi_{M}=\psi_{\mathbf{Q}}\circ\operatorname{Tr}_{M/\mathbf{Q}}$. Applying [29, Prop. 3] to both $\lambda_{\ell}$ and $\lambda_{\ell}\chi_{M,\ell}$ yields (20). As $\chi_{M,\infty}=1$, it suffices to check (21) for $v$ finite. Moreover, the characters $\lambda_{v}$ and $\chi_{M,v}$ are unramified for $v\nmid\ell D$, hence both sides of (21) are $1$. Finally, for $v$ dividing $D$, $\chi_{M,v}$ is unramified while $\lambda_{v}$ is tamely ramified and $\psi_{M_{v}}$ has conductor $1$, implying by [7, (5.5.1)] that $W_{v}(\lambda_{v}\chi_{M,v})/W_{v}(\lambda_{v})=\chi_{M,v}(\xi^{1+1\cdot 1})=1$. ∎
3.4. The Bombieri–Lang Conjecture for Picard modular surfaces
In this part we follow the general strategy of [11] by adapting it to the case where the level is not neat, the main point being to show that the irregularity of the Picard modular surfaces under consideration is at least $3$. This requires some new techniques, also providing new proofs to the cases treated in loc. cit.
We recall the index $2$ subgroup $\widetilde{K}^{\prime\prime}$ of the maximal open compact subgroup $\widetilde{K}^{\circ}$ of $\widetilde{G}(\mathbf{A}_{f})$ introduced in (13). Given $r\in\mathbf{Z}_{\geqslant 1}$ and a prime $\ell$ inert in $M$, we let $\widetilde{K}^{\prime\prime}_{0}(\ell^{r})$ be the subgroup of $\widetilde{K}^{\prime\prime}$ whose component at $\ell$ is the depth $r$ Iwahori subgroup $I_{r}$ and we let
$K^{\prime\prime}_{0}(\ell^{r})=\widetilde{K}^{\prime\prime}_{0}(\ell^{r})\cap G(\mathbf{A}_{f})\text{  and } \Gamma^{\prime\prime}_{0}(\ell^{r})=G(\mathbf{Q})\cap K^{\prime\prime}_{0}(\ell^{r})\cdot G(\mathbf{R}).$
Consider an automorphic representation $\pi\in\Pi\left(\lambda_{c}\chi_{M},{\lambda^{-1}_{c}}_{\!\!\!\!\!|M^{1}}\chi^{-2}\omega\right)$ having non-zero $K^{\prime\prime}_{0}(\ell^{r})$-invariants. By Propositions 1.3 and 1.5, for all finites place $v\neq\ell$ we must have $\pi_{v}=\pi_{n,v}$ with $\chi_{v}$ unramified, in which case it follows from (14) that the line $\bigotimes\limits_{p\mid D}\pi_{n,p}^{K^{\prime\prime}_{p}}$ is fixed by $K^{\prime\prime}_{D}$. By (18), $\pi$ contributes to $q\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)$ if either $\pi_{\ell}=\pi_{n,\ell}$ and $W(\lambda_{c}^{3}\chi^{3}_{M})=1$, or $\pi_{\ell}=\pi_{c,\ell}$ and $W(\lambda_{c}^{3}\chi^{3}_{M})=-1$. As the choice of a character $\chi_{\ell}:\mathbf{Q}_{\ell^{2}}^{1}\to\mathbf{C}^{\times}$ determines uniquely, up to an ideal class character, a global character $\chi_{M}$ unramified outside $\ell$, the irregularity formula (18) becomes (we have omitted $\omega\in\widehat{\mathscr{C}\!\ell_{M}^{1}}$ as it is trivial on $\mathbf{Q}_{\ell^{2}}^{1}$)
(22) $\tfrac{1}{h}\cdot q\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)=\sum_{\chi_{\ell},\text{ s.t.}W(\lambda_{c}^{3}\chi_{M}^{3})=1}\dim(\pi_{n,\chi_{\ell}}^{I_{r}})+\sum_{\chi_{\ell},\text{ s.t.}W(\lambda_{c}^{3}\chi_{M}^{3})=-1}\dim(\pi_{c,\chi_{\ell}}^{I_{r}}),$
where $\Pi(\lambda_{0}\chi_{M,\ell},\chi^{-2}_{\ell})=\{\pi_{n,\chi_{\ell}},\pi_{c,\chi_{\ell}}\}$ denotes the local Arthur packet attached to the quadratic unramified character $\lambda_{0}:\mathbf{Q}_{\ell^{2}}^{\times}\to\mathbf{C}^{\times}$ and to $\chi_{\ell}:\mathbf{Q}_{\ell^{2}}^{1}\to\mathbf{C}^{\times}$.
Remark 3.4.
The study of the conductor of the supercuspidal non-generic representation $\pi_{c,\chi_{\ell}}$ appears to be very delicate. Indeed, even in the depth 0 case (i.e., trivial $\chi$), some preliminary computations suggest that it does not contain non-zero $K_{T}$-invariant vectors. In a previous paper ([11]), we mistakenly switched $\pi_{c}$ and $\pi_{2}$ (which are in the same $L$-packet) in the proof of Propositions 3.6 and 3.8 in loc. cit.. It did not affect Theorem 0.2 there at all, but in Theorem 0.1 when $W(\lambda^{3})=(-1)^{d}$, we need to replace $\Gamma_{1}(\mathfrak{C})$ by $\Gamma_{1}(\mathfrak{Cq})$ for any $\mathfrak{q}$ prime of $F$ relatively prime to $\mathfrak{C}$ and non-split in $M$. We overcome the difficulty caused by $\pi_{\mathfrak{q},c}$ not having $\Gamma_{1}(\mathfrak{q})$-invariants by using an appropriate twist $\lambda_{c}\chi_{M}$ of $\lambda_{c}$ for which $W(\lambda_{c}^{3}\chi^{3}_{M})=(-1)^{d-1}$ and the corresponding $\pi_{\mathfrak{q},n}$ has the requisite invariants. Adapting this twisting method at ramified places, allows us to establish a variant of Theorem 0.3 from loc. cit.. This will be taken up elsewhere.
Henceforth, we will only use a lower bound for the irregularity corresponding to the contribution of everywhere non-tempered automorphic representations. Namely, combining (22) with Lemma 3.3, we have
(23) $q\left(Y_{\Gamma^{\prime\prime}_{0}(\ell^{r})}^{\ast}\right)\geqslant h\cdot\sum_{\chi_{\ell}}\dim(\pi_{n,\chi_{\ell}}^{I_{r}}),$
where the sums runs over characters $\chi_{\ell}:\mathbf{Q}_{\ell^{2}}^{1}\to\mathbf{C}^{\times}$ such that $\chi_{\ell}(-1)=(-1)^{a(\chi_{\ell})+(D-3)/4}$.
Proposition 3.5.
We recall that $D>3$ is odd and $h$ denotes the class number of $M=\mathbf{Q}(\sqrt{-D})$.
  1. (i)
    If $D\equiv 3\pmod{8}$ then $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(\ell)}\right)=q(Y^{\ast}_{\Gamma^{\prime\prime}})=h$ and $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(\ell^{2r})}\right)\geqslant(r+1)h$, for $r\in\mathbf{Z}_{\geqslant 1}$.
  2. (ii)
    If $D\equiv 7\pmod{8}$ then $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(27)}\right)\geqslant h$, $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(5^{6})}\right)\geqslant 3h$ and $q\left(Y^{\ast}_{\Gamma_{0}^{\prime\prime}(\ell^{3})}\right)\geqslant 3h$, for $\ell\geqslant 7$.
Proof.
(i) We can take $\chi=\mathbf{1}$ in (23). The claim follows from Proposition 1.5 and Corollary 1.8.
(ii) In this case, we use tamely ramified $\chi_{\ell}$ in (23), i.e. $a(\chi_{\ell})=1$, such that $\chi_{\ell}(-1)=1$. For $\ell\geqslant 3$, there are precisely $\frac{\ell-1}{2}$ such characters, if $3\nmid(\ell+1)$, and $\frac{\ell-5}{2}$ choices, if $3\mid(\ell+1)$. In particular, there are at least $3$ choices for all $\ell\geqslant 7$. As $K_{T}$ contains a conjugate $I_{3}$, the claim then follows from Theorem 1.9 supplemented, when $\ell=5$, by a numerical computation showing that $\pi_{n,\chi_{5}}$ has non-zero $I_{6}$-invariants for $\chi_{5}$ such that $\chi_{5}(-1)=-1$ and $a(\chi_{5})=2$. ∎
The following results from Faltings [13] (see [11, §3.2] for details).
Theorem 3.6.
Let $\widetilde{K}$ an open compact subgroup of $\widetilde{G}(\mathbf{A}_{f})$ and let $\Gamma=\widetilde{G}(\mathbf{Q})\cap\widetilde{K}\widetilde{G}(\mathbf{R})$. If $q\left(Y^{\ast}_{\Gamma}\right)\geqslant 3$, then $Y^{\ast}_{\widetilde{K}}$ satisfies the Bombieri–Lang Conjecture.
As $2$ splits in $M$ for $D\equiv 7\pmod{8}$, Proposition 3.5 has the following consequence.
Corollary 3.7.
If $D\equiv 3\pmod{8}$, then the Bombieri–Lang Conjecture holds for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(\ell^{4})}$, and even for $Y^{\ast}_{\widetilde{K}^{\prime\prime}}$, when $h\geqslant 3$. If $D\equiv 7\pmod{8}$, then the Bombieri–Lang Conjecture holds for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(\ell^{3})}$ for $\ell\geqslant 7$, and also for $Y^{\ast}_{\widetilde{K}^{\prime\prime}(3^{7})}$ and $Y^{\ast}_{\widetilde{K}^{\prime\prime}(5^{6})}$.
3.5. Proof of Main Theorem
At different stages of the proof we will remove finite sets and deal with them in the last step. By §2.1 an abelian variety $A$ as in the Theorem B defines a $k$-rational point on $Y_{\widetilde{K}^{\prime\prime}}$, where $\widetilde{K}^{\prime\prime}$ is defined in (13). If $A[\ell^{r}]$ admits a full $k$-rational flag (or equivalently, a $k$-rational isotropic line), we claim that $A$ defines a $k$-rational point on $Y_{\widetilde{K}^{\prime\prime}_{0}(\ell^{r})}$, where $\widetilde{K}^{\prime\prime}_{0}(\ell^{r})\subset\widetilde{K}^{\prime\prime}$ is the subgroup whose component at $\ell$ consists of elements whose reduction modulo $\ell^{r}$ belong to the standard Borel subgroup $\widetilde{B}(\mathbf{Z}/\ell^{r}\mathbf{Z})$ of $\widetilde{G}(\mathbf{Z}/\ell^{r}\mathbf{Z})$. Indeed, by assumption one knows that there is some Borel subgroup containing the image of $\operatorname{Gal}_{k}$ acting on $A[\ell^{r}]$. However $\widetilde{G}(\mathbf{Q}_{\ell})$ acts transitively on isotropic lines (because isometries between hermitian subspaces always extend), hence all Borel subgroups are conjugated by $\widetilde{G}(\mathbf{Q}_{\ell})$, and in fact by $\widetilde{K}_{\ell}^{\circ}=\widetilde{G}(\mathbf{Z}_{\ell})$ (using Iwasawa decomposition). As $\widetilde{K}^{\prime\prime}$ is a normal subgroup of $\widetilde{K}^{\circ}$ we deduce that the Galois image is contained in $\widetilde{K}^{\prime\prime}\cap\widetilde{K}_{0}(\ell^{r})=\widetilde{K}^{\prime\prime}_{0}(\ell^{r})$, proving the claim.
By Corollary 3.7, $Y^{\ast}_{\widetilde{K}^{\prime\prime}_{0}(\ell^{7})}$ satisfies the Strong Bombieri–Lang Conjecture. In particular all its $k$-rational points lie in a subvariety $Z$ defined over $k$ which is a finite union of points and curves.
Let us now take one of the (finitely many) geometrically connected curve $C$ in $Z$, and after removing finitely many of its points (which would not affect the wanted result), we may assume that $C$ is contained in the smooth open $U$ from §2.2. In particular, there exists a family $f:A\to C$ of abelian $3$-folds of Picard type.
As in §2.3 let $\Gamma_{C}\subset\widetilde{K}_{\ell}^{\circ}=\widetilde{G}(\mathbf{Z}_{\ell})$ be the image of the étale fundamental group acting on the $\ell$-adic Tate module of the generic fiber of the family. By Cartan’s theorem (see [30, LG5.42]), $\Gamma_{C}$ is an $\ell$-adic Lie group hence admits a Lie algebra $\mathfrak{g}_{\mathbf{Z}_{\ell}}$. By Bogomolov [2] the Lie algebra $\mathfrak{g}_{\mathbf{Q}_{\ell}}=\mathfrak{g}_{\mathbf{Z}_{\ell}}\otimes_{\mathbf{Z}_{\ell}}\mathbf{Q}_{\ell}$ is algebraic, namely it is the Lie algebra of the Zariski closure of $\Gamma_{C}$ in $\widetilde{G}(\mathbf{Q}_{\ell})$, the latter being furthermore reductive over $\mathbf{Q}_{\ell}$ by Faltings [12, Thm. 3]. By the Mumford–Tate Conjecture, which is known for is known for abelian $3$-folds (see e.g. [5]), we know that $\mathfrak{g}_{\mathbf{Q}_{\ell}}$ is the Lie algebra of the Mumford–Tate group $\operatorname{MT}(A)\otimes_{\mathbf{Q}}\mathbf{Q}_{\ell}$. As $C$ has positive dimension, it has to contain non-CM points, whose Mumford–Tate group is not abelian. By Lemma 2.3, the Lie subalgebra $\mathfrak{g}_{\mathbf{Q}_{\ell}}\cap\mathfrak{su}(3,\mathbf{Q}_{\ell})$ contains a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$. By [4, Thm. 1.1] applied to the abelian family $f:A\to C$ there exist $B>0$ such that for all $x\in C(k)$ outside a finite set $C_{\rho}$ we have
(24) $\Gamma_{C}:\Gamma_{x}]\leqslant B,$
where $\Gamma_{x}=\rho_{A_{x},\ell}(\operatorname{Gal}_{k})$ with $A_{x}$ the abelian $3$-fold of Picard type corresponding to $x$.
Lemma 3.8.
There exists $r=r(C)\in\mathbf{Z}$ such that $\Gamma_{x}\supset\exp\left(\mathfrak{su}(2,\ell^{r}\mathbf{Z}_{\ell})\right)$ for all $x\in C(k){\smallsetminus}C_{\rho}$.
Proof.
We fix an exponential map on $\mathfrak{su}(2,\mathbf{Q}_{\ell})$ so that $\Gamma_{C}\supset\exp\left(\mathfrak{su}(2,\mathbf{Z}_{\ell})\right)$. Using that a subgroup of index at most $B$ contains a normal subgroup of index at most $B!$, (24) implies that $\Gamma_{x}$ contains $B!\cdot\exp\left(\mathfrak{su}(2,\mathbf{Z}_{\ell})\right)=\exp\left(\mathfrak{su}(2,\ell^{r}\mathbf{Z}_{\ell})\right)$, where $r$ is the $\ell$-adic valuation of $B!$. ∎
It follows that for all $r\geqslant r(C)$ and for all $A$ as above corresponding to a $k$-rational point on $C{\smallsetminus}C_{\rho}$, $A[\ell^{r}]$ does not admit a full $k$-rational flag. Finally, applying Proposition 2.4 to the finitely many remaining $k$-rational points, yields an integer $r$ such that all $k$-rational points in $Y_{\widetilde{K}^{\prime\prime}_{0}(\ell^{r})}$ are of CM type, completing the proof of the Theorem.
Acknolwedgements: We would like to thank our respective institutions, and also the TIFR where part of the work was done. We would like to thank A.  Cadoret, T. Graber, M.  Goreski, B. Gross, H.  Hida, A. Jorza, K.-W. Lan, B. Mazur, J.  Nekovář, D.  Prasad, A.  Raghuram, D. Rohrlich, J.-P. Serre and J.  Tilouine for helpful discussions.